### v2

Abstract
Empirical Welfare Maximization (EWM) is a framework that can be used to select welfare program eligibility policies based on data. This paper extends EWM by allowing for uncertainty in estimating the budget needed to implement the selected policy, in addition to its welfare. Due to the additional estimation error, I show there exist no rules that achieve the highest welfare possible while satisfying a budget constraint uniformly over a wide range of DGPs. This differs from the setting without a budget constraint where uniformity is achievable. I propose an alternative trade-off rule and illustrate it with Medicaid expansion, a setting with imperfect take-up and varying program costs.
Keywords: empirical welfare maximization, heterogeneous treatment effects and costs, cost-benefit analysis, local asymptotics
JEL classification codes: C14, C44, C52
1 Introduction
When a welfare program induces varying benefits across individuals, and when resources are scarce, policymakers naturally want to prioritize eligibility to individuals who will benefit the most. Based on experimental data, cost-benefit analysis can inform policymakers on which subpopulations to prioritize, but these subpopulations might not align with any available eligibility policy such as an income threshold. Kitagawa and Tetenov (2018) propose a statistical rule, Empirical Welfare Maximization (EWM), that can directly select an eligibility policy from a set of available policies based on the experimental data. For example, if available policies take the form of income thresholds, EWM considers the problem of maximizing the expected benefits in the population
$\max_{t\leq\overline{t}}\ E[\text{benefit}\cdot\mathbf{1}\{\text{income}\leq t\}]$
and approximates the optimal threshold based on benefits estimated from experimental data. Recent papers have argued the EWM approach is attractive under a wide range of data distributions, including Athey and Wager (2021), among others. Notably the average benefits under the eligibility policy selected by EWM converges to the highest attainable level as the sample size grows. This property can be thought of as uniform asymptotic welfare efficiency.
In practice policymakers often face budget constraints, but only have imperfect information about whether a given eligibility policy satisfies the budget constraint. First, there may be imperfect take-up: eligible individuals might not participate in the welfare program, resulting in zero cost to the government, e.g. Finkelstein and Notowidigdo (2019). Second, costs incurred by eligible individuals who participate in the welfare program may vary considerably, largely driven by individuals’ different needs but also many other factors, e.g. Finkelstein et al. (2017). Both considerations are hard to predict ex ante, implying that the potential cost of providing eligibility to any given individual is unknown at the time of designing the eligibility policy. Unobservability of the potential cost requires estimation based on experimental data, contributing to uncertainty in the budget estimate of a given eligibility policy.
For this empirically relevant setting where the budget needed to implement an eligibility policy involves an unknown cost, this paper introduces a new property of statistical rules, namely asymptotic feasibility. Policies are feasible if they satisfy the budget constraint in the target population; otherwise, they are referred to as infeasible. A statistical rule is asymptotically feasible if given a large enough experimental sample, the statistical rule is very likely to select feasible eligibility policies. Since a potential cost overrun due to an underestimate of the budget can be very costly to policymakers, asymptotic feasibility is also a desirable property in the current setting, especially for policymakers who are extremely cautious about cost overruns. This paper answers three questions in the current setting with unknown cost: is there any statistical rule that achieves uniformly good performance for a wide range of data distributions as in the original EWM literature, whether the obvious extension of the existing EWM statistical rule remains attractive, and what are some alternative statistical rules.
Firstly as a novel theoretical contribution, I quantify a class of reasonable data distributions that is particularly challenging for statistical rules. Specifically, no statistical rule can be uniformly welfare-efficient and feasible simultaneously over this class of data distributions. A notable example within this class is the expansion of welfare programs, such as tax credits to incentivize labor force participation, which can “pay for themselves,” as demonstrated in Hendren and Sprung-Keyser (2020), because they have zero net cost to the government. This impossibility result provides theoretical characterization of constrained settings that need to be ruled out for any statistical rule to achieve uniformity.
Second, I show the direct extension of the existing EWM statistical rule is not appealing in the setting with unknown cost. The reason is that this sample-analog rule ignores the estimation error in the estimated budget constraint, which has non-negligible consequences even when the sample size is large. For data distributions mentioned earlier where the budget constraint is exactly binding, the welfare loss does not vanish with sample size. The probability of selecting infeasible policies also does not vanish with sample size, but a simple modification to the sample-analog rule can ensure that feasible policies are selected with high probability. These considerations are relevant in real-world settings, as demonstrated by a calibrated simulation study.
Finally I propose the trade-off rule, which can select infeasible eligibility policies, but only if borrowing money and exceeding the budget constraint are not too costly to justify the marginal gain in welfare from violating the budget constraint. I show the trade-off rule is uniformly asymptotically welfare efficient, which is an improvement over the direct extension of the EWM rule.
To illustrate the statistical rules proposed in this paper, I consider a budget-constrained Medicaid expansion. Medicaid is a government-sponsored health insurance program intended for the low-income population in the United States. Based on the Oregon Health Insurance Experiment (OHIE) conducted in 2008, I derive Medicaid expansion policies using statistical rules. Oregon’s current Medicaid expansion policy is based on a threshold for household income only. In my application, I examine whether health can be improved by allowing the income threshold to vary by the number of children in the household, setting the budget constraint to the current level. Using the sample-analog rule, the selected policy limits eligibility to be lower than the current level to meet the estimated budget constraint. However, the cost estimates are highly variable, indicating considerable uncertainty about whether the selected policy is welfare-efficient in the population. In contrast, the trade-off rule selects an eligibility policy that expands eligibility for many households above the current level, especially those with children. The higher level occurs because, based on the OHIE data, Medicaid improves health for these households, and the additional health benefit from violating the budget constraint exceeds the cost of doing so, assuming a reasonable upper bound on the monetary value on the health benefit.
The rest of the paper proceeds as follows. Section 1.1 discusses related work in more detail. Section 2 presents theoretical results. Sections 3 and 4 discuss properties of two statistical rules. Section 5 conducts a simulation study and Section 6 considers an empirical example of designing a more flexible Medicaid expansion policy for the low-income population in Oregon. Section 7 concludes. Proofs can be found in the Appendix. Supporting lemmas, additional results and computational details can be found in the Online Appendix.
1.1 Literature review
This paper is related to the traditional literature on cost-benefit analysis, e.g. Dhailiwal et al. (2013), and to the recent literature on EWM, e.g. Kitagawa and Tetenov (2018), Rai (2019), Athey and Wager (2021) and Mbakop and Tabord-Meehan (2021). More broadly, this paper contributes to a growing literature on statistical rules in econometrics, including Manski (2004), Dehejia (2005), Hirano and Porter (2009), Stoye (2009), Chamberlain (2011), Bhattacharya and Dupas (2012), Demirer et al. (2019), Yata (2021) and Kitagawa et al. (2022), among others.
The traditional cost-benefit analysis compares the cost and benefit of a given welfare program. The effect of program eligibility is first estimated based on a randomized control trial (RCT), and then converted to a monetary benefit for calculating the cost-benefit ratio. For example, Gelber et al. (2016) and Heller et al. (2017) compare the efficiency of various crime prevention programs based on their cost-benefit ratios. However, the cost-benefit ratio is only informative for whether this welfare program should be implemented with the fixed eligibility policy as implemented in the RCT.
The literature on statistical rules in econometrics has also developed a definition for optimality of statistical rules. Manski (2004) considers the minimax-regret, defined to be loss in expected welfare achieved by the statistical rule relative to the welfare achieved by the theoretically optimal eligibility policy. In the absence of any constraint, under the theoretically optimal eligibility policy, anyone with positive benefit from the welfare program would be assigned with eligibility. The minimax regret is the upper bound on the regret that results from not knowing the data distribution. Manski (2004) argues a statistical rule is preferred if its minimax-regret converges to zero with the sample size, and analyzes the minimax-regret for the class of statistical rules that only selects eligibility policy based on subsets of the observed covariates. Stoye (2009) shows that with continuous covariates and no functional form restrictions on the set of policies, minimax regret does not converge to zero with the sample size because the theoretically optimal policy can be too difficult to approximate by statistical rule. Kitagawa and Tetenov (2018) avoid this issue by imposing functional form restrictions. They propose the EWM rule, which starts with functional form restrictions on the class of available policies, and then selects the policy with the highest estimated benefit (empirical welfare) based on an RCT sample. They prove the optimality of EWM in the sense that its regret converges to zero at the minimax rate. Importantly, the regret is defined to be loss in expected welfare relative to the maximum achievable welfare in the constrained class, which avoids the negative results of Stoye (2009). Athey and Wager (2021) propose doubly-robust estimation of the average benefit, which leads to an optimal rule even with quasi-experimental data. Mbakop and Tabord-Meehan (2021) propose a Penalized Welfare Maximization assignment rule which relaxes restrictions of the policy class.
The existing EWM literature has not addressed budget constraints with an unknown cost. Kitagawa and Tetenov (2018) consider a capacity constraint, which they enforce using random rationing. Random rationing is not ideal as it uses the limited resource less efficiently than accounting for the cost of providing the welfare program to an individual. When there is no restriction on the functional form of the eligibility policy, Bhattacharya and Dupas (2012) demonstrate that given a capacity constraint, the optimal eligibility policy is based on a threshold on the benefit of the welfare program to an individual. When the cost of providing the welfare program to an individual is heterogeneous, however, budget constraints can be more complicated than capacity constraints, and require estimation. Carneiro et al. (2020) considers the optimal choice of covariate collection in order to maximize the precision of the estimation for the average treatment effect. While they also consider a constrained decision problem, the budget constraint can be verified directly. They also allow for a more complicate trade-off between additional covariates and additional observations. Sun et al. (2021) propose a framework for estimating the optimal rule under a budget constraint when there is no functional form restriction. The main contribution of this paper is to characterize theoretical properties of statistical rules when allowing both functional form restrictions and budget constraints with an unknown cost.
As explained in a prior version of this paper (Sun, 2021), the current setting of EWM with constraint shares the same mathematical structure as another important setting: fairness constraints across sensitive subgroups. Viviano and Bradic (2024) cast welfare of sensitive subgroups as multiple objective functions for policymakers, and solve the constrained optimal policy via the Pareto frontier. Kock and Preinerstorfer (2024) consider a penalized objective function that penalizes violations to the constraint. The trade-off rule proposed in this paper takes a similar form by penalizing budget overrun, but is tailored to linear objectives and constraints.
2 Theoretical results
2.1 Motivation and setup
I begin by setting up a general constrained optimization problem, which depends on the following random attributes of an individual:
$A=(\Gamma,R,X)\in\mathcal{A}\subseteq\mathbb{R}^{2+p}.$ (1)
Here $\Gamma$ is benefit from the treatment for the individual, $R$ is the cost to the policymaker of providing the individual with the treatment, and $X\in\mathcal{X}\subset\mathbb{R}^{p}$ denotes their $p$-dimensional characteristics. The individual belongs to a population that can be characterized by the joint distribution $P$ on the random attributes $A$. The unknown distribution $P\in\mathcal{P}$ is from a class of distributions $\mathcal{P}$.
A policy $g(X)\in\{0,1\}$ determines the treatment status for an individual with observed characteristics $X$, where 1 is treatment and 0 is no treatment. Let $\mathcal{G}$ denote the class of policies policymakers can choose from. The optimization problem is to find a policy with maximal benefit while subject to a constraint on its cost: [^1] Following Kitagawa and Tetenov (2018), I implicitly assume the maximizer exists in $\mathcal{G}$ with the notation in (2).
$\max_{g\in\mathcal{G}}E_{P}[\Gamma\cdot g(X)]\ s.t.\ E_{P}[R\cdot g(X)]$ $\leq k.$ (2)
If the policymaker does not have a fixed budget but still want to account for cost, the scalar $\Gamma$ can be the difference in benefit and cost. Since a fixed budget is common, I impose a harsh constraint.
The benefit-cost attributes $(\Gamma,R)$ of any given individual may or may not be observed. When the benefit-cost attributes are observed, policymakers observe i.i.d. $A_{i}=(\Gamma_{i},R_{i},X_{i})$ of individual $i$ in a random sample of the target population. When the benefit-cost attributes are unobserved, the focus of this paper is the setting where policymakers can construct their estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$ in a random sample of sample size $n$ along with the characteristics $X_{i}$ from an experiment or quasi-experiment that satisfy Assumption 3.1, as discussed later.
Applying the Law of Iterated Expectation, the constrained optimization problem (2) can be written as
$\max_{g\in\mathcal{G}}E_{P}[E_{P}[\Gamma\mid X]\cdot g(X)]\ s.t.\ E_{P}[E_{P}[R\mid X]\cdot g(X)]\leq k.$ (3)
When the eligibility policy can be based on any characteristics whatsoever, the class of available policies is unrestricted i.e. $\mathcal{G}=2^{\mathcal{X}}$. In this unrestricted class, when the cost is non-negative, the above expression makes clear that the optimal eligibility policy is based on thresholding by the benefit-cost ratio $E_{P}[\Gamma\mid X]/E_{P}[R\mid X]$ where the numerator and the denominator are respectively the average effects conditional on the observed characteristics (CATE) and the conditional average resource required. Online Appendix B provides a formal statement.
Given a random sample, to approximate the optimal eligibility policy $g_{P}^{\ast}$, one can estimate the benefit-cost ratio based on the estimated CATE and the estimated conditional average resource required. The resulting statistical rule selects eligibility policies that are thresholds based on the estimated benefit-cost ratio. The challenge is that the selected eligibility policy can be hard to implement when the estimated benefit-cost ratio is a complicated function of $X$. Restrictions on the policy class $\mathcal{G}$ address this issue. A common restriction is to consider thresholds based directly on $X$, e.g. assigning eligibility when an individual’s income is below a certain value.
Restrictions on the policy class $\mathcal{G}$ mean that there might not be closed-form solutions to the population problem (2). In particular, the constrained optimal eligibility policy $g_{P}^{\ast}$ might not be an explicit function of the CATE and the conditional average resource required. Therefore it might be difficult to directly approximate $g_{P}^{\ast}$ based on the estimated CATE and the estimated conditional average resource required. However, this is not an obstacle to deriving guarantees for the statistical rules. As I demonstrate later, the derivation does not require the knowledge of the functional form of the constrained optimal eligibility policy $g_{P}^{\ast}$.
I next specialize the constrained optimization problem to selecting eligibility policy for welfare programs with the example of Medicaid expansion. In the example of implementing welfare programs, policies take the form of eligibility policies. I restrict attention to non-randomized policies as in the leading example of welfare programs, deterministic policies such as income thresholds are more relevant. While the optimization problem policymakers face in practice can take on different forms, constrained optimization problems of this mathematical structural are ubiquitous in social science. Theoretically oriented readers may proceed directly to Section 2.2.
Example 2.1.
Welfare program eligibility policy under budget constraint
Suppose the government wants to implement some welfare program. The treatment in this example is eligibility for such welfare program. Due to a limited budget, the government cannot make eligibility universal and can only provide eligibility to a subpopulation. To use the budget efficiently, policymakers consider the constrained optimization problem (2). In this example, the policy $g(X)$ assigns an individual to eligibility based on their observed characteristics $X$, and is usually referred to as an eligibility policy. I denote $\Gamma$ to be the benefit experienced by an individual after receiving eligibility for the welfare program. Specifically, let $(Y_{1},Y_{0})$ denote the potential outcomes that would have been observed if an individual were assigned with and without eligibility, respectively. The benefit from eligibility policy is therefore defined as $\Gamma\coloneqq Y_{1}-Y_{0}$. Note that maximizing benefit is equivalent to maximizing the outcomes (welfare) under the utilitarian social welfare function: $E_{P}[Y_{1}\cdot g(X)+Y_{0}(1-g(X))]$. I denote $R$ to be the potential cost from providing an individual with eligibility for the welfare program. Both $\Gamma$ and $R$ are unobserved at the time of assignment and will need to be estimated.
Policymakers might be interested in multiple outcomes for an in-kind transfer program. Hendren and Sprung-Keyser (2020) capture benefits by the willingness to pay (WTP). Assuming eligible individuals make optimal choices across multiple outcomes, the envelope theorem allows policymakers to focus on benefit in terms of one particular outcome.
Medicaid Expansion
Medicaid is a government-sponsored health insurance program intended for the low-income population in the United States. Up till 2011, many states provided Medicaid eligibility to able-bodied adults with income up to 100% of the federal poverty level. The 2011 Affordable Care Act (ACA) provided resources for states to expand Medicaid eligibility for all adults with income up to 138% of the federal poverty level starting in 2014.
Suppose policymakers want to maximize the health benefit of Medicaid by adopting a more flexible expansion policy. Specifically, they relax the uniform income threshold of 138% and allows the income thresholds to vary with the number of children in the household. It may be infeasible to implement a new expansion policy if it costs more than the current one. Therefore I impose the constraint that the more flexible policy must cost no more than the current one.
Correspondingly, the constrained optimization problem (2) sets $\Gamma$ to be the health benefit from Medicaid, $R$ to be the excess per-enrollee cost of Medicaid relative to the current expansion policy, and the appropriate threshold to be $k=0$. The characteristics $X$ include both income and number of children in the household.
The policy class in this example includes income thresholds that can vary with the number of children in the household:
$\mathcal{G}=\left\{g(x)=\begin{cases}\mathbf{1}\{\text{income}\leq\beta_{0}\}&,\ \text{numchild}=0\\ \vdots\\ \mathbf{1}\{\text{income}\leq\beta_{j}\}&,\ \text{numchild}=j\end{cases}\right\}$ (4)
for characteristics $x=(\text{income},\ \text{numchild})$ and $\ \beta_{j}\geq 0$.
2.2 Desirable properties for statistical rules
To simplify the notation, I define the welfare function and the budget function:
$W(g;P)=E_{P}[\Gamma\cdot g(X)],\ B(g;P)=E_{P}[R\cdot g(X)].$ (5)
and the constrained optimal policy $g_{P}^{\ast}$ is therefore the solution to
$\max_{g\in\mathcal{G}}W(g;P)\ s.t.\ B(g;P)$ $\leq k.$
As explained in Example 2.1 from Section 2, under a utilitarian social welfare function, maximizing the benefit with respect to eligibility policy is equivalent to maximizing the welfare, which is why I refer to $W(g;P)$ as the welfare function. The welfare function and the budget function are both deterministic functions from $\mathcal{G}\rightarrow\mathbb{R}$. The index by the distribution $P$ highlights that welfare and budget of policy $g$ vary with $P$, and in particular, whether a policy $g$ satisfies the budget constraint depends on which distribution $P$ is of interest.
When the benefit-cost attributes $(\Gamma,R)$ are unobserved and the distribution $P$ is unknown, both the welfare function and the budget function are unknown functions. Denote by $\widehat{g}$ a statistical rule that selects an eligibility policy after observing some experimental data of sample size $n$ distributed according to $P^{n}$. This section provides formal definitions for two desirable properties of $\widehat{g}$.
Definition 2.1.
A statistical rule $\widehat{g}$ is pointwise asymptotically welfare-efficient under the data distribution $P$ if for any $\epsilon>0$
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0,$ (6)
and uniformly asymptotically welfare-efficient over the class of distributions $\mathcal{P}$ if for any $\epsilon>0$
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{W(\widehat{g};P)-W(g_{P}^{\ast};P)<-\epsilon\right\}=0.$
A statistical rule is pointwise asymptotically feasible under the data distribution $P$ if
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\{B(\widehat{g};P)>k\}=0,$
and uniformly asymptotically feasible over the class of distributions $\mathcal{P}$ if
$\lim_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{B(\widehat{g};P)>k\}=0.$ (7)
$\blacksquare$
The above two properties build on the existing EWM literature. For the first property, the current EWM literature evaluates statistical rules by whether they attain at least $W(g_{P}^{\ast};P)$ in expectation over repeated sample draws as $n\rightarrow\infty$. Instead, I focus on convergence in probability, where the probability of $\widehat{g}$ selecting eligibility policies that achieve strictly lower welfare than $g_{P}^{\ast}$ approaches zero as $n\rightarrow\infty$. In the setting of the existing EWM literature, the constrained optimal policy $g_{P}^{\ast}$ is also the unconstrained optimal policy $\mathcal{G}$. Therefore it is impossible for any statistical rule $\widehat{g}$ to select a policy that achieves higher value than $g_{P}^{\ast}$. In my setting, however, the constrained optimal policy $g_{P}^{\ast}$ is not necessarily the unconstrained optimal policy. Thus, I allow the statistical rule $\widehat{g}$ to select a policy that achieves higher welfare than $g_{P}^{\ast}$ for all data distributions, albeit at the cost of violating the budget constraint.
The second property is new to the EWM literature. It imposes that given a large enough sample size, the statistical rule $\widehat{g}$ is unlikely to select infeasible eligibility policies that violate the budget constraint, so that it is “asymptotically feasible”. Asymptotic feasibility of statistical rules is specific to the current setting where the budget constraint involves unknown cost. Exactly satisfying a fixed budget constraint without the smallest violation is the most conservative way to articulate policy makers’ preferences.
While both are desirable properties, the next section shows a negative result that it is impossible for a statistical rule to satisfy both properties when the data distribution is unknown and belongs to a sufficiently rich class of distributions $\mathcal{P}$.
2.3 An impossibility result
Similar to the classic “impossibility results” (Pfanzagl, 1998; Hirano and Porter, 2012), one may conjecture that there cannot exist any statistical rule that approximates the constrained optimum $g^{\ast}_{P}$ uniformly well in the neighbourhood of $P_{0}$, where $g^{\ast}_{P_{h_{n}}}$ can differ arbitrarily from $g^{\ast}_{P_{0}}$ even though the underlying distributions $P_{h_{n}}$ are arbitrarily close to $P_{0}$. This section formalizes this impossibility result that no statistical rule can be both uniformly asymptotically welfare-efficient and uniformly asymptotically feasible in a sufficiently rich class of distributions $\mathcal{P}$. The point of discontinuity $P_{0}$ is characterized in Assumptions 2.2 and 2.3.
Assumption 2.1.
Contiguity. There exists a distribution $P_{0}\in\mathcal{P}$ under which a non-empty set of eligibility policies satisfying the constraint exactly $\mathcal{G}_{0}=\left\{g:B(g;P_{0})=k\right\}$. Furthermore, the class of distributions $\mathcal{P}$ includes a sequence of data distributions $\{P_{h_{n}}\}$ contiguous to $P_{0}$, under which for all $g\in\mathcal{G}_{0}$, there exists some $C>0$ such that
$\sqrt{n}\cdot\left(B(g;P_{h_{n}})-k\right)>C.$
Assumption 2.1 assumes there exists a sequence of data distributions $\{P_{h_{n}}\}$ that differ only marginally relative to $P_{0}$. Moreover, the budget function $B(g;P_{h_{n}})$, evaluated at polices that meet the budget constraint exactly under $P_{0}$, approaches $B(g;P_{0})$ from above. In Online Appendix B.2, I give more primitive assumptions under which Assumption 2.1 is guaranteed to hold, requiring the sequence to be differentiable in quadratic mean at $P_{0}$, and along the sequence, the budget function $B(g;P_{h_{n}})$ is twice continuously differentiable at $P_{0}$ with positive derivatives. These primitive assumptions are relatively weak, and have also been considered in the literature to construct the local parametrization around $P_{0}$, e.g. Hirano and Porter (2009) .
Assumption 2.2.
Binding constraint. Under the data distribution $P_{0}$, the constraint is satisfied exactly at the constrained optimum i.e. $B(g_{P_{0}}^{\ast};P_{0})=k$.
Assumption 2.3.
Separation. Under the data distribution $P_{0}$, $\exists\epsilon>0$ such that for any feasible policy $g$, whenever
$\left|B(g;P_{0})-B(g_{P_{0}}^{\ast};P_{0})\right|>0,$
we have
$W(g_{P_{0}}^{\ast};P_{0})-W(g;P_{0})>\epsilon.$
Equivalently, $W(g_{P_{0}}^{\ast};P_{0})$ is separated from that of other feasible policies with different $B(g;P_{0})$.
Consider a one-dimensional policy class $\mathcal{G}$, e.g. income thresholds. Figure 2.1 illustrates a distribution $P_{0}$ that satisfies both Assumptions 2.2 and 2.3, while both the welfare function $W(g;P_{0})$ and the budget function $B(g;P_{0})$ are still continuous in $g$, satisfying Assumption 2.1. Importantly, the constrained optimal policy $g_{P_{0}}^{\ast}$ satisfies the budget constraint exactly, i.e. $B(g_{P_{0}}^{\ast};P_{0})=k$, but is separated from the rest of feasible eligibility policies such that there exists a neighborhood around $g_{P_{0}}^{\ast}$ where feasible policies can achieve welfare gain without any effect on the budget.
Theorem 2.1.
Suppose Assumption 2.1 holds for the class of data distributions $\mathcal{P}$. For $P_{0}\in\mathcal{P}$ considered in Assumption 2.1, suppose it also satisfies Assumptions 2.2 and 2.3. Then no statistical rule can be both uniformly asymptotically welfare-efficient and uniformly asymptotically feasible. In particular, if a statistical rule $\widehat{g}$ is pointwise asymptotically welfare-efficient and pointwise asymptotically feasible under $P_{0}$, then it is not uniformly asymptotically feasible.
Figure 2.1 provides some intuition for Theorem 2.1. Note that if a statistical rule $\widehat{g}$ is pointwise asymptotically welfare-efficient and pointwise asymptotically feasible under $P_{0}$, then it has to select eligibility policies close to $g_{P_{0}}^{\ast}$ with high probability over repeated sample draws distributed according to $P_{0}^{n}$ as $n\rightarrow\infty$.
Under Assumption 2.1, the class of distributions $\mathcal{P}$ is sufficiently rich so that along a sequence of data distributions $\{P_{h_{n}}\}$ that is contiguous to $P_{0}$ as $n\rightarrow\infty$, the budget functions $B(g;P_{h_{n}})$ converge to $B(g;P_{0})$ while $B(g_{P_{0}}^{\ast};P_{h_{n}})>k$, i.e. $g_{P_{0}}^{\ast}$ is not feasible under $P_{h_{n}}$. Figure 2.1 showcases $P_{1}$ as one distribution from this sequence. The contiguity between $\{P_{h_{n}}\}$ and $P_{0}$ implies that the statistical rule $\widehat{g}$ must select policies close to $g_{P_{0}}^{\ast}$ with high probability under $P_{h_{n}}^{n}$ as well. However, the policy $g_{P_{0}}^{\ast}$ is infeasible under $P_{h_{n}}$, and therefore the statistical rule $\widehat{g}$ cannot be asymptotically feasible under $P_{h_{n}}$.
Figure 2.1: Illustration for Assumptions 2.1-2.3 underlying Theorem 2.1
Notes: This figure plots welfare functions $W(g;P)$ and budget functions $B(g;P)$ for populations distributed according to $P_{0}$ (blue solid lines) or $P_{1}$ (red dashed lines), where $P_{1}$ is a distribution from the sequence of distributions $\{P_{h_{n}}\}$ that is contiguous to $P_{0}$ under Assumption 2.1. The distribution $P_{0}$ satisfies Assumptions 2.2 and 2.3. The $x$-axis indexes a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$ for a one-dimensional continuous characteristic $X_{i}$ with support on $[0,1]$, e.g. eligibility policies based on income thresholds. The black dotted line marks the budget threshold $k$. The bold blue dot marks $g_{P_{0}}^{\ast}$, the constrained optimal eligibility policy under $P_{0}$. The bold red dot marks $g_{P_{1}}^{\ast}$, the constrained optimal eligibility policy under $P_{1}$.
The impossibility result is relevant in real-world settings. It is possible to encounter situations where an expansion of a welfare program has no impact on the budget but leads to a strictly positive increase in welfare, as described in Assumptions 2.2 and 2.3. While the nominal cost of welfare program eligibility may be positive, the Marginal Value of Public Funds (MVPF) framework emphasizes that the relevant cost is the net cost, which incorporates fiscal externalities, such as increased tax revenue if the program increases individuals’ incomes. Specifically, Hendren and Sprung-Keyser (2020) estimate fourteen welfare programs (out of 133) to have negative or zero net cost to the government, which implies these programs “pay for themselves”, aligning with Assumptions 2.2 and 2.3. As a result, the impossibility result suggests that statistical rules designed for such cases may be highly sensitive to small changes in the underlying distributions even in large samples.
3 Sample-analog rule
The previous negative result implies that no statistical rule can be both uniformly asymptotically welfare-efficient and uniformly asymptotically feasible. Thus, policymakers might want to consider statistical rules that satisfy one of these two properties. A direct extension to the existing approach in the EWM literature, the sample-analog rule, might be a natural candidate. In this section, I show the direct extension is neither uniformly asymptotically welfare-efficient nor uniformly asymptotically feasible.
I first describe the EWM approach (Kitagawa and Tetenov, 2018) and its direct extension. Since $(\Gamma,R)$ involves potential outcomes, they are often unobserved and require estimation based on RCT that introduces estimation errors in addition to sampling errors. Section 3.2 describes how to construct individuals’ benefit and cost estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$. To highlight the drawback of the direct extension to EWM, I first consider settings where we observe an experimental data of sample size $n$ where $(\Gamma,R)$ is directly observable, i.e. $(\Gamma_{i}^{\ast},R_{i}^{\ast})=(\Gamma_{i},R_{i})$. The goal of the simplification is to highlight that the non-uniformity I show below can arise from sampling errors alone.
One can estimate the welfare function and the budget function using their sample-analog versions:
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i}).$ (8)
A direct extension to the existing approach in the EWM literature is a statistical rule that solves the sample version of the population constrained optimization problem (2):
$\widehat{g}_{\text{sample}}\in\arg\max_{\widehat{B}_{n}(g)\leq k}\ \widehat{W}_{n}(g).$ (9)
The subscript “sample” emphasizes how this approach verifies whether a policy satisfies the constraint by comparing the sample analog $\widehat{B}_{n}(g)$ with $k$ directly, i.e. imposes a sample-analog constraint. If no policy satisfies the constraint, then I set $\widehat{g}_{\text{sample}}$ to not assign any eligibility, i.e. $\widehat{g}_{\text{sample}}(x)=0$ for all $x\in\mathcal{X}$.
3.1 Welfare inefficiency of the sample-analog rule
A key insight from Kitagawa and Tetenov (2018) is that without a constraint, the sample-analog rule is uniformly asymptotically welfare-efficient. Unfortunately this intuition does not extend to the current setting where the constraint involves an unknown cost. There are common data distributions under which the amount of welfare loss does not vanish even as the sample size gets larger. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which is based on thresholds of a one-dimensional continuous characteristic $X$. Suppose the policymakers know benefit is positive for everyone so that welfare function is strictly increasing, and only need to estimate whether a given threshold satisfies a capacity constraint due to imperfect take-up. Furthermore, suppose the experiment sample observes take-up $R_{i}$ so that the only uncertainty arises from sampling errors. Proposition 3.1 shows settings with some zero take-ups satisfy Assumptions 2.2 and 2.3, and the sample-analog rule is neither pointwise asymptotically welfare efficient nor pointwise asymptotically feasible.
Proposition 3.1.
One-dimensional threshold and imperfect take-up. Consider the special case where under distribution $P$, the benefit $\Gamma>0$ almost surely and the cost $R\in\{0,1\}$ is binary. Suppose $R=0$ for $X\in[\underline{t},\overline{t}]$, but $Pr_{P}\{R=1\mid X\}\in(0,1)$ otherwise. Then $P$ also satisfies Assumptions 2.2 and 2.3. As $n\rightarrow\infty$, the sample-analog rule $\widehat{g}_{\text{sample}}$ would violate the budget constraint half the time, and incur at least $\epsilon>0$ loss in welfare the other half the time:
$Pr_{P^{n}}\left\{W(g_{P}^{*};P)-W(\widehat{g}_{\text{sample}};P)\geq\epsilon\right\}\rightarrow 50\%.$
Figure 3.1: Illustration for Proposition 3.1
(a) infeasible
(b) suboptimal
Notes: This figure plots the budget function $B(g;P)$ (blue dotted line) and its sample-analogs $\widehat{B}_{n}(g)$ (black wriggly line) based on two different observed samples in panel (a) and (b) respectively. The $x$-axis indexes a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$ for a one-dimensional continuous characteristic $X_{i}$ with support on $[0,1]$. The black dotted line marks the budget threshold $k$. The constrained optimal threshold $g_{P}^{\ast}$ is $t=0.5$. The sample-analog rule $\widehat{g}_{\text{sample}}$ selects an infeasible threshold in panel (a) and selects a suboptimal threshold in panel (b).
Figure 3.1 illustrates the setup of Proposition 3.1, where the sampling uncertainty can be particularly problematic for the sample-analog rule. Since policymakers know the benefit is positive for everyone, $\widehat{g}_{\text{sample}}$ takes a simple form of the highest threshold where the sample-analog constraint is satisfied exactly. The driving force behind the failure of $\widehat{g}_{\text{sample}}$ as described in Proposition 3.1 is that due to sampling uncertainty, whether a policy satisfies the sample-analog constraint is an imperfect measure of whether it satisfies the constraint in the population.
Since the sample-analog rule $\widehat{g}_{\text{sample}}$ restricts attention to policies that satisfy the sample-analog constraint, there is no guarantee the selected policy is actually feasible. This is very likely to happen when there is welfare gain in exceeding the budget constraint as in the setup of Proposition 3.1 where $W(g;P)$ is strictly increasing in $g$. Therefore, the sample-analog rule $\widehat{g}_{\text{sample}}$ is not asymptotically feasible under $P$. As illustrated in Figure 3.1, after observing a sample depicted in panel (a), the sample-analog rule picks an infeasible threshold because the sample-analog constraint is still satisfied there. However, as argued later in Corollary 3.1, the probability of large budget violations vanishes as the sample size increases.
The more problematic case is illustrated in Figure 3.1 panel (b), where the sample-analog rule picks a suboptimal threshold because the sample-analog constraint is violated at the constrained optimum $g_{P}^{\ast}$. In the setup of Proposition 3.1, when the sample-analog rule $\widehat{g}_{\text{sample}}$ misses $g_{P}^{\ast}$, it is guaranteed to select a suboptimal policy and therefore it is welfare-inefficient under $P$ even asymptotically. This is because the welfare function strictly increases when the budget function remains constant in the neighborhood of the budget constraint. I demonstrate that the same issue extends to broader contexts by employing a simulation calibrated to a real-world DGP from the OHIE in Section 5.
3.2 Estimates for benefit and cost
In the remainder of the paper, I present constructive results that modify the sample-analog rule and propose a new rule. To lay the groundwork, this section outlines the benefit and cost estimates. The appropriate expressions for these estimates depend on the type of observed data. Below I state the estimates formed based an RCT that randomly assigns the eligibility, which is the leading case of Kitagawa and Tetenov (2018). The observed data $\{A_{i}^{\ast}\}_{i=1}^{n}$ consists of i.i.d. observations $A_{i}^{\ast}=(Y_{i},Z_{i},D_{i},X_{i})\in\mathcal{A}^{\ast}$. The distribution of $A_{i}^{\ast}$ is induced by the distribution of $(Y_{1},Y_{0},R,X)$ as in the population, as well as the sampling design of the RCT. Here $D_{i}$ is an indicator for being in the eligibility arm of the RCT, $Y_{i}$ is the observed outcome and $Z_{i}$ is the observed cost of providing eligibility to an individual participating in the RCT. The observed cost is mechanically zero if an individual is not randomized into the eligibility arm. The estimates for $(\Gamma,R)$ are
$\Gamma_{i}^{\ast}=\alpha(X_{i},D_{i})\cdot Y_{i},\ R_{i}^{\ast}=\frac{D_{i}}{p(X_{i})}\cdot Z_{i},$ (10)
where $\alpha(X_{i},D_{i})=\frac{D_{i}}{p(X_{i})}-\frac{1-D_{i}}{1-p(X_{i})}$ and $p(X_{i})$ is the propensity score, the probability of receiving eligibility conditional on the observed characteristics. Since the sampling design of an RCT is known, the propensity score is a known function of the observed characteristics.
Assumption 3.1.
Estimation quality. The recentered empirical processes $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$ defined in (8) converge to mean-zero Gaussian processes $G_{P}^{W}$ and $G_{P}^{B}$ uniformly over $g\in\mathcal{G}$, with covariance functions $\Sigma_{P}^{W}(\cdot,\cdot)$ and $\Sigma_{P}^{B}(\cdot,\cdot)$ respectively:
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i})-W(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{W}$
$\left\{\sqrt{n}\cdot\left(\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i})-B(g;P)\right)\right\}_{g\in\mathcal{G}}$ $\rightarrow_{d}G_{P}^{B}$
Moreover, the convergence holds uniformly over $P\in\mathcal{P}$. The covariance functions are uniformly bounded, with diagonal entries bounded away from zero uniformly over $g\in\mathcal{G}$. There is a uniformly consistent estimator $\widehat{\Sigma}^{B}(\cdot,\cdot)$ of the covariance function $\Sigma_{P}^{B}(\cdot,\cdot)$.
Online Appendix B.3 gives primitive assumptions under which Assumption 3.1 is guaranteed to hold for $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$ constructed using an RCT such as in (10) or an observational study, assuming unconfoundedness and strong overlap. As standard in the literature, I need to restrict the complexity of the policy class $\mathcal{G}$. The policy class of income thresholds considered in this paper is in fact a rather simple class with VC-dimension $d+1$ where $d$ is the number of different thresholds as in (4).
3.3 Modification that ensures feasibility
Let $\hat{\mathcal{G}}=\{g\in\mathcal{G}:\widehat{B}_{n}(g)\leq k\}$ denote the set of policies that $\widehat{g}_{\text{sample}}$ can choose from, which contains policies that do not violate the sample-analog of the budget constraint. Unsurprisingly, for any finite sample, $\hat{\mathcal{G}}$ can always contain infeasible policies, sometimes of sizable budget violations. The probability that $\hat{\mathcal{G}}$ contains a policy that violates the population budget constraint by a fixed amount $c$ is the follows:
$Pr_{P^{n}}\left\{\exists g:g\in\hat{\mathcal{G}}\ and\ B(g;P)>k+c\right\}$ (11)
The corollary stated below shows the chance that a large amount of budget violation of $c>0$ occurs is smaller when there is less variability in $\widehat{B}_{n}(g)$. The chance also vanishes to zero as the sample size gets larger.
Corollary 3.1.
Under Assumption 3.1, an upper bound of the probability (11)can be approximated by the CDF of $\underset{g\in\mathcal{G}}{\inf}\frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}$ evaluated at $\frac{-\sqrt{n}\cdot c}{\max_{g\in\mathcal{G}}\Sigma^{B}(g,g)^{1/2}}$.
A simple modification to the sample-analog rule can reduce the probability of selecting infeasible policies. Instead of $\hat{\mathcal{G}}$, let $\widehat{g}_{\alpha}$ choose policies from a subset $\hat{\mathcal{G}}_{\alpha}$ defined in Theorem 3.1 and maximize $\widehat{W}_{n}(g)$ as before. Then as a direct consequence of Theorem 3.1, with probability at least $1-\alpha$ this modified rule is guaranteed to not mistakenly choose infeasible eligibility policies.
Theorem 3.1.
Suppose Assumption 3.1 holds for the class of data distributions $\mathcal{P}$. Collect eligibility policies in
$\hat{\mathcal{G}}_{\alpha}=\left\{g:g\in\mathcal{G}\text{ and }\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\right\},$ (12)
where $c_{\alpha}$ is the $\alpha$-quantile from $\underset{g\in\mathcal{G}}{\inf}\frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}$ for $G_{P}^{B}$ the Gaussian process defined in Assumption 3.1, and $\widehat{\Sigma}^{B}(\cdot,\cdot)$ the consistent estimator for its covariance function. Then
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\hat{\mathcal{G}}_{\alpha}\cap\mathcal{G}^{+}\neq\emptyset\}<\alpha,$ (13)
where $\mathcal{G}^{+}=\{g:B(g;P)>k\}$ is the set of infeasible eligibility policies.
Note that in (12) the sample-analog constraint is tightened by $c_{\alpha}\cdot\frac{\widehat{\Sigma}^{B}(g,g)^{1/2}}{\sqrt{n}}$ where $c_{\alpha}$ is negative, which means the class $\hat{\mathcal{G}}_{\alpha}$ only includes eligibility policies where the constraint is slack in the sample. The sample-analog constraint is tightened proportionally to the standard error to reflect that $\widehat{B}_{n}(g)$ might be particularly noisy for some $g$. The tightening therefore shrinks inversely proportional to the (square root of) sample size because intuitively larger sample size reduces the sampling uncertainty.
In practice, $\alpha$ may be set at the conventional level, e.g. 5% for an explicit coverage guarantee. The drawback is that $\widehat{g}_{\alpha}$ can be welfare inefficient. One may therefore let $\alpha_{n}\rightarrow 0$ as sample size increases to achieve both pointwise asymptotically welfare-efficient and pointwise asymptotically feasible for distributions where the constraint is slack at the constrained optimal eligibility policy. Corollary 3.2 formalizes the rate requirement below.
Corollary 3.2.
If the data distribution $P$ satisfies Assumption 3.1 and induces a constrained optimal policy at which the constraint is slack, then the modified sample-analog rule
$\widehat{g}_{\alpha_{n}}\in\arg\max_{g\in\hat{\mathcal{G}}_{\alpha_{n}}}\widehat{W}_{n}(g)$ (14)
is asymptotically welfare-efficient and asymptotically feasible under $P$ for $\alpha_{n}\rightarrow 0$ at a rate such that $c_{\alpha_{n}}=o(n^{1/2})$.
4 Trade-off rule
Exceeding the budget constraint in the population might be desirable if it achieves higher welfare and does not cost too much. How to model such trade-off depends on the specific welfare program. Section 4.1 first explores possible forms of the trade-off that are of practical relevance, and then focuses on the trade-off where the marginal cost per unit of violating the constraint has a known upper bound. The upper bound is context specific and as an example, Section 6 discusses how to set it for Medicaid expansion. Relaxing the constraint results in weakly higher welfare than that of the constrained optimal policy $g_{P}^{\ast}$. Section 4.2 derives a statistical rule that implements such trade-off in the sample, and is shown to be uniformly asymptotically welfare-efficient.
4.1 Form of the trade-off
Note that the original constrained optimization problem (2) may be reformulated as
$\max_{g\in\mathcal{G}}\inf_{\lambda\geq 0}\ W(g;P)-\lambda\cdot(B(g;P)-k).$ (15)
where $\lambda$ measures the marginal gain of relaxing the constraint. In many settings where the benefit is monetary, it may be natural to think for every dollar spent right above the budget constraint, there is a welfare gain of $\lambda$ dollars.
This formulation implies that policymakers are willing to enforce the constraint at all costs since $\lambda$ is unbounded. This formulation may not reflect the real objective, however, as policymakers might be willing to trade off violations of the constraint against gains in welfare only to a certain extent, bounding $\lambda\in[0,\overline{\lambda}]$ and resulting in a new objective function
$\max_{g\in\mathcal{G}}\min_{\lambda\in[0,\overline{\lambda}]}\ W(g;P)-\lambda\cdot(B(g;P)-k).$ (16)
Using the notation $(x)_{+}$ to denote the positive part of $x\in\mathbb{R}$, the inner optimization problem of (16) can be written as
$\min_{\lambda\in[0,\overline{\lambda}]}\ W(g;P)-\lambda\cdot(B(g;P)-k)=W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}.$
Note, however, this reformulated objective does not involve optimization over $\lambda$: the solution to (16) is equivalent to the solution to a non-smooth but piecewise linear optimization problem:
$\widetilde{g}_{P}\in\arg\max_{g\in\mathcal{G}}\ W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}.$ (17)
Note that this trade-off problem (17) is equivalent to an objective function faced by policymakers when they can finance the deficit $(B(g;P)-k)_{+}$ through borrowing at a fixed interest rate $\overline{\lambda}$, after converting welfare into monetary terms. Therefore, the solution $\widetilde{g}_{P}$ can relax the budget constraint and achieves weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. The next lemma formalizes this observation.
Lemma 4.1.
For any $\overline{\lambda}\in[0,\infty)$, the solution to the trade-off problem (17) achieves weakly higher welfare than the constrained optimum: $W(\widetilde{g}_{P};P)\geq W(\widetilde{g}_{P};P)-\overline{\lambda}\cdot(B(\widetilde{g}_{P};P)-k)_{+}\geq W(g_{P}^{\ast};P)$ for any distribution $P$. Moreover, the violation to the budget constraint is upper bounded by $\frac{W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$.
4.2 Welfare efficiency of the trade-off rule
Given a new objective function (17) that trades off the gain and the cost from violating the constraint, the goal is to derive a statistical rule that is likely to select eligibility policies that maximize the new objective function. Consider the trade-off statistical rule defined as
$\widehat{g}_{\text{tradeoff}}\in\arg\max_{g\in\mathcal{G}}\ \widehat{W}_{n}(g)-\overline{\lambda}\cdot(\widehat{B}_{n}(g)-k)_{+},$ (18)
where the subscript “tradeoff” highlights that this statistical rule is able to relax the constraint by trading off the gain and the cost from violating the constraint. Since $\widehat{g}_{\text{tradeoff}}$ solves a sample-analog version of (17), I show it consistently achieves the same value of (17) under weak conditions. Theorem 4.1 further verifies that it consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$. I leave to future research to verify whether the trade-off rule is minimax rate optimal.
Theorem 4.1.
Suppose Assumption 3.1 holds for the class of data distributions $\mathcal{P}$. Then the trade-off rule $\widehat{g}_{\text{tradeoff}}$ defined in (18) consistently achieves welfare that is weakly higher than $W(g_{P}^{\ast};P)$ for all $P\in\mathcal{P}$ and is uniformly asymptotically welfare-efficient under $\mathcal{P}$. Moreover, with probability approaching one, the violation to the budget constraint is upper bounded by $\frac{W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$ uniformly over $\mathcal{P}$.
To gain intuition for the above results, I note that under Assumption 3.1, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ consistently achieves the same value of (17) as $\widetilde{g}_{P}$ over repeated sample draws as $n\rightarrow\infty$ uniformly over $\mathcal{P}$. Furthermore, Lemma 4.1 shows the trade-off solution $\widetilde{g}_{P}$ achieves weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$ for any data distribution $P$. Therefore, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ is asymptotically welfare-efficient uniformly over $\mathcal{P}$. At the same time, larger trade-off coefficient $\bar{\lambda}$ implies smaller violation to the budget constraint, relative to the welfare gain $W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)$.
5 Monte Carlo Simulations
To ensure the practical relevance of the simulation, I calibrate to the distribution of the data from the OHIE, which is also used for empirical illustration in Section 6, based on Example 2.1 of Section 2. The benefit is defined to be the increase in the probability of reporting good subjective health after receiving Medicaid eligibility, and the cost is defined to be health care expenditure that needs to be reimbursed by Medicaid, in excess of the current policy. I defer the details on the construction of the estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$ to Section 6.1. Table 6.1 presents the basic summary statistics for the estimates $(\Gamma_{i}^{\ast},R_{i}^{\ast})$. For the purpose of this simulation study, the OHIE represents the population $P$, and I take these estimates as the true benefit and cost $(\Gamma,R)$. Under this simulation design, I can solve for the constrained optimal policy as
$g_{P}^{\ast}\in\arg\max_{g\in\mathcal{G},\ B(g;P)\leq 0}W(g;P)$
where $\left(W(g;P),B(g;P)\right)$ are the sample analogs in the OHIE sample. The policy class $\mathcal{G}$ includes income thresholds that can vary with the number of children. Under the current definition of the cost, a policy is feasible if it incurs a negative cost. Table 6.1 shows the policy of expanding eligibility to the full population, while increasing health by 4%, violates the budget constraint by \$358 per enrollee. The constrained optimal policy therefore only makes a share of the population eligible. The maximum feasible welfare is given by $W(g_{P}^{\ast};P)=3.8\%$, an increase of 3.8% in reporting good subjective health. The cost associated with the constrained optimal policy is $B(g_{P}^{\ast};P)=-\$0.6$, meaning per enrollee $g_{P}^{\ast}$ costs \$0.6 less than the current policy.
5.1 Simulation results
Table 5.1 compares the performance of various statistical rules $\widehat{g}$ through 500 Monte Carlo iterations. At each iteration, I randomly draw observations from the OHIE sample to form a random sample. I simulate with the same sample size as the original sample to hold the amount of sampling uncertainty constant. Given the random sample, I collect eligibility policies chosen by each of the following statistical rules:
  • sample-analog rule $\widehat{g}_{\text{sample}}$,
  • modified sample-analog rule $\widehat{g}_{\alpha}$ described in Theorem 3.1 with $\alpha$ set to 5%,
  • trade-off rule $\widehat{g}_{\text{tradeoff}}$ with trade-off coefficient $\bar{\lambda}$ described in Section 6.2.
I evaluate the welfare function and the budget function $\left(W(g;P),B(g;P)\right)$ for a given policy in the original OHIE sample. Averages over 500 iterations provide simulation evidence on the properties of the above statistical rules, as shown in Table 5.1.
Table 5.1: Simulation results
Statistical rule sample-analog modified trade-off
$\widehat{g}_{\text{sample}}$ $\widehat{g}_{\alpha}$ $\widehat{g}_{\text{tradeoff}}$
Prob. of selecting infeasible policies 35.4% 8% 79.6%
Prob. of selecting suboptimal policies 87.0% 98.6% 37.6%
Average welfare loss 0.06 0.60 -0.02
Average cost -\$3 -\$57 \$105
Notes: This table reports properties of statistical rules $\widehat{g}$, as averaged over 500 simulations. Row 1 reports the probability that the rule selects an eligibility policy that violates the budget constraint, i.e. $Pr_{P^{n}}\{B(\widehat{g};P)>0\}$. Row 2 reports the probability that the rule achieves strictly less welfare than the constrained optimal policy $g_{P}^{\ast}$, i.e. $Pr_{P^{n}}\{W(\widehat{g};P)<W(g_{P}^{\ast};P)\}$. Row 3 reports the average welfare loss of the rule relative to the maximum feasible welfare, i.e. $\frac{E_{P^{n}}\left[W(g_{P}^{\ast};P)-W(\widehat{g};P)\right]}{W(g_{P}^{\ast};P)}$. Row 4 reports the average cost of the policies selected by the rule, i.e. $E_{P^{n}}\left[B(\widehat{g};P)\right]$.
Row 1 of Table 5.1 illustrates that it is possible for all three statistical rule $\widehat{g}$ to select infeasible policies. A lower probability of selecting infeasible policies suggests the rule is closer to achieving asymptotic feasibility. In the distribution calibrated to the OHIE sample, the original sample-analog rule $\widehat{g}_{\text{sample}}$ might not be asymptotically feasible as it can select infeasible eligibility policies in 35.4% of the draws. In contrast, Theorem 3.1 guarantees that a simple modification $\widehat{g}_{\alpha}$ selects infeasible eligibility policies in less than 5% of the draws, regardless of the distribution. Simulation confirms such guarantee as the mistakes only happen 8% of the time.
Row 2 of Table 5.1 illustrates that it is possible for all three statistical rule $\widehat{g}$ to achieve weakly higher welfare than the constrained optimal policy $g_{P}^{\ast}$. This can happen when $\widehat{g}$ selects an infeasible policy. A lower probability of selecting suboptimal policies suggests the rule is closer to achieving asymptotic welfare-efficiency. Theorem 4.1 implies that the trade-off rule $\widehat{g}_{\text{tradeoff}}$ is uniformly asymptotically welfare efficient while there is no such guarantee for the sample-analog rule $\widehat{g}_{\text{sample}}$.
In the distribution calibrated to the OHIE, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ on average achieves higher welfare than the sample-analog rule $\widehat{g}_{\text{sample}}$. As shown in row 3 of Table 5.1, the welfare loss of $\widehat{g}_{\text{tradeoff}}$ is -2% of the maximum feasible welfare $W(g_{P}^{\ast};P)$, compared to 6% for $\widehat{g}_{\text{sample}}$. However, its improvement can be at the cost of violating the budget constraint more often than $\widehat{g}_{\text{sample}}$, at a rate of 79.6%. Though as shown in row 4 of Table 5.1, on average the violation is limited, which confirms Theorem 4.1.
6 Medicaid expansion: empirical illustration
Example 2.1 of Section 2 explains a more flexible Medicaid expansion policy that would allow the income thresholds to vary with the number of children. This section illustrates this example using data from the the Oregon Medicaid Health Insurance Experiment (OHIE).
6.1 Data
I use the experimental data from the OHIE, where Medicaid eligibility ($D_{i}$) was randomized in 2007 among Oregon residents who were low-income adults, but previously ineligible for Medicaid, and who expressed interest in participating in the experiment. Finkelstein et al. (2012) include a detailed description of the experiment and an assessment of the average effects of Medicaid on health and health care utilization. I include a cursory explanation here for completeness.
The original OHIE sample consists of 74,922 individuals (representing 66,385 households). Of these, 26,423 individuals responded to the initial mail survey, which collects information on income as percentage of the federal poverty level and number of children, which are the characteristics $X_{i}$ of interest for targeting. [^2] More accurately, I follow Sacarny et al. (2020) to approximate number of children by the number of family members under age 19 living in house as reported on the initial mail survey. I exclude individuals who did not respond to the initial survey from my sample, which differs from the sample analyzed in Finkelstein et al. (2012) as I focus on individuals who responded both to the initial and the main surveys from the OHIE. Due to this difference, the expansion policies selected using my sample do not directly carry their properties to the population underlying the original OHIE sample, as the distributions of $X$ differ. After one year, the main survey collects data related to health ($Y_{i}$), health care utilization ($C_{i}$) and actual enrollment in Medicaid ($M_{i}$), which allows me to construct estimates for the benefit and cost of Medicaid eligibility $(\Gamma,R)$. Therefore I further exclude individuals who did not respond to the main survey from my sample.
For health ($Y_{i}$), I follow the binary measurement in Finkelstein et al. (2012) based on self-reported health, where an answer of “poor/fair” is coded as $Y_{i}=0$ and “excellent/very good/good” is coded as $Y_{i}=1$. For health care utilization ($C_{i}$), the study collected measures of utilization of prescription drugs, outpatient visits, ER visits, and inpatient hospital visits. Finkelstein et al. (2012) annualize these utilization measures to turn these into spending estimates, weighting each type by its average cost (expenditures) among low-income publicly insured non-elderly adults in the Medical Expenditure Survey (MEPS). Note that health and health care utilization are not measured at the same scale, which requires rescaling when I consider the trade-off between the two. I address this issue in Section 6.2. Lastly, since the enrollment in Medicaid still requires an application, not everyone eligible in the OHIE eventually enrolled in Medicaid, which implies $M_{i}\leq D_{i}$.
Given the setup of the OHIE, Medicaid eligibility ($D_{i}$) is random conditional on household size (number of adults in the household) entered on the lottery sign-up form and survey wave. While the original experimental setup would ensure randomization given household size, the OHIE had to adjust randomization for later waves of survey respondents (see the Appendix of Finkelstein et al. (2012) for more details). Denote the confounders (household size and survey wave) with $V_{i}$, and define the propensity score as $p(V_{i})=Pr\{D_{i}=1\mid V_{i}\}$. If the propensity score is known, then the construction of the estimates follows directly from the formula (10). However, the adjustment for later survey waves means I need to estimate the propensity score, and I adapt the formula (10) following Athey and Wager (2021) to account for the estimated propensity score.
Specifically, define the conditional expectation function (CEF) of a random variable $U_{i}$ as $\gamma^{U}=E[U_{i}\mid V_{i},D_{i}$]. Since $V_{i}$ in my case is discrete, I use a fully saturated model to estimate the propensity score $\widehat{p}(V_{i})$ and the CEF $\widehat{\gamma}^{U}(V_{i},D_{i})$. I then form the estimated Horvitz-Thompson weight with the estimated propensity score as $\widehat{\alpha}(V_{i},D_{i})=\frac{D_{i}}{\widehat{p}(V_{i})}-\frac{1-D_{i}}{1-\widehat{p}(V_{i})}$. For health benefit due to Medicaid eligibility, define the estimate $\Gamma_{i}^{\ast}=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$. The 2014 Medicaid spending was roughly \$6,000 per adult enrollee in Oregon, according to the expenditure information obtained from MACPAC (2019). To formalize the budget constraint that the per enrollee cost of the proposed policy cannot exceed the 2014 policy, I need to account for imperfect take-up because not everyone eligible for Medicaid would enroll. For the per enrollee excess cost relative to the current level, define the estimate $R_{i}^{\ast}=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(W_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$ where $Z_{i}=C_{i}-\$6,000\cdot M_{i}$.
Table 6.1: Summary statistics of the OHIE sample by number of children
Number of children Sample size Sample mean of $\Gamma_{i}^{\ast}$ Sample mean of $R_{i}^{\ast}$
0 5,758 3.1% \$651
(0.01) (138)
1 1,736 10.3% \$348
(0.02) (247)
$\geq 2$ 2,641 1.5% -\$275
(0.02) (150)
Full sample 10,135 4.0% \$358
(0.01) (97)
Notes: This table presents summary statistics on the sample of individuals who responded to both the initial and the main surveys from the Oregon Health Insurance Experiment (the OHIE sample). The first three rows represent individuals living with different number of children (family members under age 19), and the last row is the aggregate. The estimate for benefit $\Gamma_{i}^{\ast}$ is an estimate for the increase in the probability of an individual reporting “excellent/very good/good” on self-reported health (as opposed to “poor/fair”) after receiving Medicaid eligibility. The estimate for cost $R_{i}^{\ast}$ is an estimate for individual’s health care expenditure that needs to be reimbursed by Medicaid, in excess of the current policy. Standard errors are shown in parentheses below.
Table 6.1 presents the summary statistics. While Online Appendix B.3 argues the estimation errors in $\Gamma_{i}^{\ast}$ and $R_{i}^{\ast}$ are asymptotically negligible, in finite samples, the cost estimates are highly variable, resulting in noisy estimate $\widehat{B}_{n}(g)$.
6.2 Budget-constrained Medicaid expansion
Figure 6.1 summarizes the selected expansion policies, which are income thresholds specific to the number of children. The sample-analog rule $\widehat{g}_{\text{sample}}$ chooses to restrict Medicaid eligibility, especially lowering the income threshold for childless individuals far below the current level. As a result, the budget estimate for the selected policy is -\$0.6, just below the threshold of zero. However, due to the large variation in the cost estimates as illustrated in Table 6.1 , meeting the sample budget constraint still involves uncertainty about whether the selected policy is welfare-efficient in the population as argued in Proposition 3.1.
To construct the trade-off rule $\widehat{g}_{\text{tradeoff}}$ as proposed in Section 4, I need to choose $\overline{\lambda}$, the upper bound on the marginal gain from violating the constraint. In my empirical illustration, the budget constraint is in terms of monetary value. The objective function, however, is measured based on self-reported health, which does not directly translate to a monetary value. Following Finkelstein et al. (2019), I convert self-reported health into value of a statistical life year (VSLY) based on existing estimates. Specifically, a conservative measure for the increase in quality-adjusted life year (QALY) when self-reported health increases from “poor/fair” to “excellent/very good/good” is roughly 0.6. The “consensus” estimate for the VSLY for one unit of QALY from Cutler (2004) is \$100,000 for the general US population. Taken these estimates together, I set $\overline{\lambda}=1/(0.6\cdot 10^{5})$. Given this upper bound, the trade-off rule $\widehat{g}_{\text{tradeoff}}$ chooses to assign Medicaid eligibility to more individuals, and to raise the income thresholds above the current level for those with children. Therefore the budget estimate for the selected policy is \$110, slightly above the budget constraint. The higher level occurs because on average the benefit estimates are positive, and the trade-off rule finds that the additional health benefit from violating the budget constraint exceeds the cost of doing so, especially those with children. While a traditional cost-benefit analysis would also recommend prioritizing those with children based on Table 6.1, the trade-off rule offers a more interpretable recommendation by directly selecting income thresholds.
Figure 6.1: More flexible Medicaid expansion policies
(a) policy selected by the sample-analog rule
(b) policy selected by the trade-off rule
Notes: This figure plots the more flexible Medicaid expansion policies selected by statistical rules based on results from the OHIE. The horizontal dashed line marks the income thresholds under the current expansion policy, which is 138% regardless of the number of children in a household. The horizontal solid lines mark the more flexible policy selected by various statistical rules, i.e. income thresholds that can vary with number of children. For each number of children, I also plot the underlying income distribution to visualize individuals below the thresholds. Panel (a) plots the policy selected by the sample-analog rule $\widehat{g}_{\text{sample}}$. Panel (b) plots the policy selected by the trade-off rule $\widehat{g}_{\text{tradeoff}}$.
7 Conclusion
In this paper, I focus on properties of statistical rules when the cost of implementing any given policy needs to be estimated. The existing EWM rule selects an eligibility policy that maximizes a sample analog of the social welfare function, and only accounts for constraints that can be verified with certainty in the population. However, in some cases, the cost of providing eligibility to any given individual might be unknown ex-ante due to imperfect take-up and heterogeneity. Therefore, in addition to asymptotic welfare-efficiency that has been studied by the EWM literature, I introduce a new desirable property of statistical rules in the setting of unknown cost, namely asymptotic feasibility, which requires the selected policy to satisfy a budget constraint when the sample size is large enough. Unlike the setting of known cost, I prove an impossibility result that no statistical rule can be uniformly asymptotically welfare-efficient and feasible. The direct extension to the existing EWM approach is no longer asymptotically welfare efficient nor asymptotically feasible for certain real-world relevant data distributions. As an alternative, I propose the trade-off rule that guarantees asymptotic welfare efficiency while ensuring any budget violations are bounded above by welfare gains. I illustrate the theoretical results using experimental data from the OHIE. A promising avenue for future research is to verify whether the trade-off rule maintains minimax rate optimality in the constrained setting, just as how the EWM rule is in the unconstrained setting.
Appendix A Proofs of theorems
Proof.
Proof of Theorem 2.1. The population is a probability space $(\Omega,\mathcal{A},P)$, which induces the sampling distribution $P^{n}$ that governs the observed sample. A statistical rule $\widehat{g}$ is a mapping $\widehat{g}(\cdot):\Omega\rightarrow\mathcal{G}$ that selects a policy from the policy class $\mathcal{G}$ based on the observed sample. Note that the selected policy $\widehat{g}(\omega)$ is still deterministic because the policy class $\mathcal{G}$ is restricted to be deterministic policies. When no confusion arises, I drop the reference to event $\omega$ for notational simplicity.
Suppose $\widehat{g}$ is asymptotically welfare-efficient and asymptotically feasible under $P_{0}$. We want to prove there is non-vanishing chance that $\widehat{g}$ selects policies that are infeasible some other distributions:
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{\omega:B(\widehat{g}(\omega);P)>k\right\}$ $>0.$ (19)
Firstly, asymptotical welfare-efficiency of $\widehat{g}$ under $P_{0}$ implies for any $\epsilon>0$ we have
$\limsup_{n\rightarrow\infty}Pr_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})>\epsilon\right\}=0.$ (20)
Asymptotical feasibility under $P_{0}$ implies $Pr_{P_{0}^{n}}\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\}\rightarrow 1.$
Consider the event $\omega^{\prime}$ where $\left|W(\widehat{g}(\omega^{\prime});P)-W(g_{P}^{\ast};P)\right|<\epsilon$ and $\widehat{g}(\omega^{\prime})$ is feasible. Asymptotic welfare-efficiency and asymptotic feasibility imply $Pr_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\rightarrow 1.$ To see this, note the probability of such event has an asymptotic lower bound of one
$Pr_{P_{0}^{n}}\left\{\omega^{\prime}:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega^{\prime});P_{0})\leq\epsilon\ and\ B(\widehat{g}(\omega^{\prime});P_{0})\leq k\right\}$
$\geq$ $Pr_{P_{0}^{n}}\left\{\omega:W(g_{p_{0}}^{\ast}P_{0})-W(\widehat{g}(\omega);P_{0})\leq\epsilon\right\}+Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})\leq k\right\}-1$
where the first two terms converge to one as $n\rightarrow\infty$ respectively under asymptotic welfare-efficiency and asymptotic feasibility.
Now consider $P_{0}$ that satisfies Assumptions 2.2 and 2.3. For $\epsilon$ in Assumption 2.3, we have $B(\widehat{g};P_{0})=B(g_{P_{0}}^{\ast};P_{0})=k$ under the event $\omega^{\prime}$, since the constraint is exactly satisfied at $g_{P_{0}}^{\ast}$ under Assumption 2.2. By Law of Total Probability, we have
$Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\geq Pr_{P_{0}^{n}}\left\{\omega^{\prime}\right\}\cdot Pr_{P_{0}^{n}}\left\{B(\widehat{g};P_{0})=k\mid\omega^{\prime}\right\}$
Then the above argument shows $Pr_{P_{0}^{n}}\left\{\omega:B(\widehat{g}(\omega);P_{0})=k\right\}\rightarrow 1$ as $n\rightarrow\infty$.
Following the notation in Assumption 2.1, denote the set of policies where the constraints bind exactly under the limit distribution $P_{0}$ by
$\mathcal{G}_{0}=\{g\in\mathcal{G}:B(g;P_{0})=k\}.$ (21)
Under Assumption 2.1, the sequence $P_{h_{n}}^{n}$ is contiguous with respect to the sequence $P_{0}^{n}$ , which means $P_{0}^{n}(A_{n})\rightarrow 0$ implies $P_{h_{n}}^{n}(A_{n})\rightarrow 0$ for every sequence of measurable sets $A_{n}$ on $\mathcal{A}^{n}$. Then $Pr_{P_{0}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\rightarrow 1$ implies there exists an $N(u)$ such that for all $n\geq N(u)$, we have $Pr_{P_{h_{n}}^{n}}\left\{A_{n}:B(\widehat{g}(A_{n});P_{0})=k\right\}\geq 1-u.$ That is, with high probability, the statistical rule $\widehat{g}$ selects policies from $\mathcal{G}_{0}$ based on the observed sample distributed according to $P_{h_{n}}^{n}$. Recall Assumption 2.1 implies for all $g\in\mathcal{G}_{0}$, for any sample size $n$, we have $B(g;P_{h_{n}})-k>C/\sqrt{n}.$ Thus this statistical rule cannot uniformly satisfy the constraint since with sample size $n\geq N(u)$, we have
$\sup_{P\in\mathcal{P}}Pr_{P^{n}}\left\{B(\widehat{g};P)>k\right\}\geq Pr_{P_{h_{n}}^{n}}\left\{B(\widehat{g};P_{h_{n}})>k\right\}\geq 1-u$ (22)
Proof.
Proof of Proposition 3.1. Consider a one-dimensional policy class $\mathcal{G}=\{g:g(x)=\mathbf{1}\{x\leq t\}\}$, which includes thresholds for a one-dimensional continuous characteristic $X$.
Suppose the budget constraint takes the form of a capacity constraint, involving a binary take-up decision $R$ (that may or may not be independent of $X$). By assumption, the probability an individual takes up the treatment for $X\in[\underline{t},\overline{t}]$ is zero but between zero and one otherwise. Then the budget function $B(t;P)\coloneqq E_{P}[R\cdot\mathbf{1}\{X\leq t\}]$ is flat in the interval $[\underline{t},\overline{t}]$ but also strictly increasing otherwise. Since $\Gamma>0$ almost surely, the welfare function $W(t;P)\coloneqq E_{P}[\Gamma\cdot\mathbf{1}\{X\leq t\}]$ is strictly increasing in $t$ and the constrained optimal policy satisfies Assumption 2.2 and is binding. Let $\inf_{B(g;P)<k}W(g_{P}^{\ast};P)-W(g;P)=E_{P}[\Gamma\cdot\mathbf{1}\{\underline{t}\leq X\leq\overline{t}\}]=\epsilon$ be the smallest amount of welfare loss from missing the binding solution $g_{P}^{\ast}$. Then $\epsilon>0$ and satisfies Assumption 2.3.
The population problem is
$\max_{t}\ W(t;P)\ s.t.\ B(t;P)\leq k,$
where $B(t;P)=k$ for $t\in[\underline{t},\overline{t}]$ by assumption. The constrained optimal threshold is therefore the highest threshold where the constraint is satisfied exactly i.e. $t^{\ast}=\overline{t}.$ This also implies $R\cdot\mathbf{1}\{X\leq t\}\sim Bernoulli(k)$ for $t\in[\underline{t},\overline{t}]$.
The sample-analog rule solves the following sample problem
$\max_{t}\frac{1}{n}\sum_{i}\Gamma_{i}\cdot\mathbf{1}\{X_{i}\leq t\}\ s.t.\ \widehat{B}_{n}(t)\leq k$
for $\widehat{B}_{n}(t)\coloneqq\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq t\}$. Given that $\Gamma>0$ almost surely, the sample-analog rule equivalently solves $\max_{t}\widehat{B}_{n}(t)\ s.t.\ \widehat{B}_{n}(t)\leq k$. However, the solution is not unique because $\widehat{B}_{n}(t)$ is a step function. To be conservative, let the sample-analog rule be the smallest possible threshold to maximize $\widehat{B}_{n}(t)$:
$\widehat{t}=\min\left\{\arg\max_{t}\{\widehat{B}_{n}(t)\ s.t.\ \widehat{B}_{n}(t)\leq k\}\right\}.$
Note that we can also write $\widehat{B}_{n}(t)=\frac{1}{n}\sum_{R_{i}=1}\mathbf{1}\{X_{i}\leq t\}$, which makes it clear that $\widehat{t}$ corresponds to ranking $X_{i}$ among individuals with $R_{i}=1$, and then picking the lowest threshold such that we assign treatment to the first $\left\lfloor k\cdot n\right\rfloor$ individuals. This also means if in the sample few individuals take up the treatment such that $\frac{1}{n}\sum_{i}R_{i}\leq k$, we can have a sample-analog rule that treats everyone up to $\max_{R_{i}=1}X_{i}$. Taken together both scenarios, we note the sample-analog rule implies the treated share in the sample is equal to
$\widehat{B}_{n}(\widehat{t})\coloneqq\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\widehat{t}\}=\min\left\{\frac{1}{n}\sum_{i}R_{i},\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}.$
Note that
$\begin{cases}\widehat{t}>\overline{t}\Leftrightarrow&B(\widehat{t};P)>B(\overline{t};P)\\ \widehat{t}<\underline{t}\Leftrightarrow&B(\widehat{t};P)<B(\underline{t};P)\Leftrightarrow W(\widehat{t};P)<W(\underline{t};P)\end{cases}$
which means whenever $\widehat{t}>\overline{t}$, the sample-analog rule violates the constraint in the population as $B(\overline{t};P)=k$; whenever $\widehat{t}<\underline{t}$, the sample-analog rule achieves strictly less welfare than $t^{\ast}$ in the population because $W(\underline{t};P)$ is strictly less than $W(t^{\ast};P)$. We next derive the limit probability for these two events. Applying Law of Total Probability, we have
$Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\widehat{B}_{n}(\widehat{t})\right\}$
$=$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}$
$+Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{1}{n}\sum_{i}R_{i}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}\geq\frac{1}{n}\sum_{i}R_{i}\right\}$ (23)
$\geq$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\ and\ \frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}$
$\geq$ $Pr_{P^{n}}\left\{\widehat{B}_{n}(\overline{t})<\frac{\left\lfloor k\cdot n\right\rfloor}{n}\right\}+Pr_{P^{n}}\left\{\frac{\left\lfloor k\cdot n\right\rfloor}{n}<\frac{1}{n}\sum_{i}R_{i}\right\}-1$ (24)
For the first term in (24), we have the following lower bound
$Pr_{P^{n}}\left\{\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}\leq\frac{k\cdot n-1}{n}\right\}$
$=$ $Pr_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}R_{i}\cdot\mathbf{1}\{X_{i}\leq\overline{t}\}-k\right)\leq-\frac{1}{\sqrt{n}}\right\}\rightarrow 0.5$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $-\frac{1}{\sqrt{n}}$ converges to zero. Denote $p_{R}=Pr\{R=1\}$. For the second term in (24), we have the following lower bound
$Pr_{P^{n}}\left\{\frac{1}{n}\sum_{i}R_{i}\geq\frac{k\cdot n}{n}\right\}$
$=$ $Pr_{P^{n}}\left\{\sqrt{n}\left(\frac{1}{n}\sum_{i}R_{i}-p_{R}\right)\geq\sqrt{n}\cdot\left(k-p_{R}\right)\right\}\rightarrow 1$
To see the convergence, we apply the Central Limit Theorem to the LHS, and note that $\sqrt{n}\cdot\left(k-p_{R}\right)$ diverges to $\infty$- ∞ for $p_{R}>k$. We thus conclude
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}\geq 0.5$
which proves $\widehat{t}$ is not pointwise asymptotically feasible under the distribution $P$.
Similar argument shows $Pr_{P^{n}}\left\{\widehat{B}_{n}(\underline{t})>\widehat{B}_{n}(\widehat{t})\right\}$ has a limit of at least one half. We thus conclude
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)<B(\underline{t};P)\right\}\geq 0.5\Leftrightarrow\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}\geq 0.5$
which proves that $\widehat{t}$ is not pointwise asymptotically welfare-efficient under the distribution $P$. Since $B(\underline{t};P)=B(\overline{t};P)$, we actually have
$\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{B(\widehat{t};P)>B(\overline{t};P)\right\}=\lim_{n\rightarrow\infty}Pr_{P^{n}}\left\{W(\widehat{t};P)<W(\underline{t};P)\right\}=0.5$
Proof.
Proof of Corollary 3.1. The event in (11) is equivalent to the event that $\hat{\mathcal{G}}$ includes at least one criteria that violates the budget constraint by $c$. The derivation for the bound therefore is based on such an event:
$Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\leq 0\right\}$
$=$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\left\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}+\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}\leq 0\right\}$
$\leq$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(B(g;P)-k\right)}{\Sigma^{B}(g,g)^{1/2}}\right\}$
$\leq$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq-\frac{\min_{B(g;P)>k+c}\sqrt{n}\left(B(g;P)-k\right)}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $Pr_{P^{n}}\left\{\min_{B(g;P)>k+c}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{B(g;P)>k+c}\Sigma^{B}(g,g)^{1/2}}\right\}$
$<$ $Pr_{P^{n}}\left\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\Sigma^{B}(g,g)^{1/2}}\leq\frac{-\sqrt{n}\cdot c}{\max_{g\in\mathcal{G}}\Sigma^{B}(g,g)^{1/2}}\right\}$
Under Assumption 3.1, the empirical process $\left\{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)\right\}$ converges to a Gaussian process $G_{P}^{B}$ for $G_{P}^{B}(\cdot)\sim\mathcal{GP}(0,\Sigma_{P}^{B}(\cdot,\cdot))$ and we have a consistent covariance estimate $\widehat{\Sigma}^{B}(\cdot,\cdot)$, which proves the corollary.
Proof.
Proof of Theorem 3.1. By construction, the limit probability for any policy in $\hat{\mathcal{G}}_{\alpha}$ to violate the budget constraint is
$Pr_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha}\ and\ B(g;P)>k\}=Pr_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $Pr_{P^{n}}\{\min_{B(g;P)>k}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$\leq$ $Pr_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
Under Assumption 3.1, uniformly over $P\in\mathcal{P}$, the empirical process $\left\{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)\right\}$ converges to a Gaussian process $G_{P}^{B}$ for $G_{P}^{B}(\cdot)\sim\mathcal{GP}(0,\Sigma_{P}^{B}(\cdot,\cdot))$ and we have a consistent covariance estimate $\widehat{\Sigma}^{B}(\cdot,\cdot)$. Then by the definition of $c_{\alpha}$, we have
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\min_{g\in\mathcal{G}}\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-B(g;P)\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}$
$=$ $\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\inf_{\text{ }g\in\mathcal{G}}\frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}\leq c_{\alpha}\}=\alpha.$
Proof.
Proof of Corollary 3.2. The first part follows from Theorem 3.1, replacing $\alpha$ with $\alpha_{n}$ in its proof. Specifically, by construction we have
$\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{B(\hat{g}_{\alpha_{n}};P)>k\}\leq\limsup_{n\rightarrow\infty}\sup_{P\in\mathcal{P}}Pr_{P^{n}}\{\exists g:g\in\hat{\mathcal{G}}_{\alpha_{n}}\ and\ B(g;P)>k\}\leq\alpha_{n}.$
For the second part, we decompose the welfare loss into
$W(g_{P}^{\ast};P)-W(\hat{g}_{\alpha_{n}};P)$
$=$ $W(g_{P}^{\ast};P)-\widehat{W}_{n}(\hat{g}_{\alpha_{n}})+\widehat{W}_{n}(\hat{g}_{\alpha_{n}})-W(\hat{g}_{\alpha_{n}};P)$
$\leq$ $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|+\widehat{W}_{n}(g_{P}^{\ast})-\widehat{W}_{n}(\hat{g}_{\alpha_{n}})$
Under the event that $g_{P}^{\ast}\in\hat{\mathcal{G}}_{\alpha_{n}}$ , we are guaranteed that the last term is non-positive. This happens with probability
$Pr_{P^{n}}\{\frac{\sqrt{n}\left(\widehat{B}_{n}(g_{P}^{\ast})-k\right)}{\widehat{\Sigma}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\leq c_{\alpha_{n}}\}\stackrel{{\scriptstyle a}}{{\sim}}\Phi\left(c_{\alpha_{n}}+\frac{\sqrt{n}\left(k-B(g_{P}^{\ast};P)\right)}{\Sigma_{P}^{B}(g_{P}^{\ast},g_{P}^{\ast})^{1/2}}\right)\rightarrow 1$
for $B(g_{P}^{\ast};P)$ strictly below the threshold. Since $c_{\alpha_{n}}$ diverges to $\infty$- ∞ at a rate slower $\sqrt{n}$, the term in the parenthesis diverges to $\infty$ as $n\rightarrow\infty$. We then apply Lemma B.4 and B.6 in the Online Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1), which shows $\sup_{g\in\mathcal{G}}|W(g;P)-\widehat{W}_{n}(g)|$ converges to zero in probability. We then conclude $\hat{g}_{\alpha_{n}}$ is asymptotically welfare-efficient under distribution $P$. ∎
Proof.
Proof of Lemma 4.1. By definition
$W(g_{P}^{\ast};P)=$ $\max_{B(g;P)\leq k}W(g;P)=\max_{B(g;P)\leq k}W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$
$\leq\max_{g\in\mathcal{G}}W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$
$=W(\widetilde{g}_{P};P)-\overline{\lambda}\cdot(B(\widetilde{g}_{P};P)-k)_{+}\leq W(\widetilde{g}_{P};P)$
and suppose $B(\widetilde{g}_{P};P)>k$, we have the following upper bound for violations to the budget constraint
$B(\widetilde{g}_{P};P)-k\leq\frac{W(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}.$
Proof.
Proof of Theorem 4.1. We first prove $\widehat{g}_{\text{tradeoff}}$ is asymptotically welfare-efficient. Denote the new objective function with $V(g;P)=W(g;P)-\overline{\lambda}\cdot(B(g;P)-k)_{+}$, whose maximizer is $\widetilde{g}_{P}$. Denote $\widehat{V}_{n}(g)=\widehat{W}_{n}(g)-\overline{\lambda}\cdot(\widehat{B}_{n}(g)-k)_{+}$ to be the sample-analog, whose maximizer is $\widehat{g}_{\text{tradeoff}}$. Apply uniform deviation bound to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $\widetilde{g}_{P}$, we have
$V(\widetilde{g}_{P};P)-V(\widehat{g}_{\text{tradeoff}};P)$
$=$ $V(\widetilde{g}_{P};P)-\widehat{V}_{n}(\widehat{g}_{\text{tradeoff}})+\widehat{V}_{n}(\widehat{g}_{\text{tradeoff}})-V(\widehat{g}_{\text{tradeoff}};P)$
$\leq$ $V(\widetilde{g}_{P};P)-\widehat{V}_{n}(\widetilde{g}_{P})+\widehat{V}_{n}(\widehat{g}_{\text{tradeoff}})-V(\widehat{g}_{\text{tradeoff}};P)$
$\leq$ $2\sup_{g}\left|V(g;P)-\widehat{V}_{n}(g)\right|$
$=$ $2\sup_{g}\left|W(g;P)-\widehat{W}_{n}(g)-\overline{\lambda}\cdot(\max\{B(g;P)-k,0\}-\max\{\widehat{B}_{n}(g)-k,0\})\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\overline{\lambda}\cdot\sup_{g}\left|\max\{\widehat{B}_{n}(g)-k,0\}-\max\{B(g;P)-k,0\}\right|$
$\leq$ $2\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|+2\overline{\lambda}\cdot\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$
The last line uses the fact that $\left|\max\{a,0\}-\max\{b,0\}\right|\leq\left|a-b\right|$. Both terms, $\sup_{g}\left|\widehat{W}_{n}(g)-W(g;P)\right|$ and $\sup_{g}\left|\widehat{B}_{n}(g)-B(g;P)\right|$ converge to zero in probability under Assumption 3.1. Specifically, Lemma B.4 and B.6 in the Online Appendix (proved in Section B.3 under primitive conditions that lead to Assumption 3.1) imply the uniform convergence in probability
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{W}_{n}(g)-W(g;P)\right|\rightarrow_{p}0,\ \sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\widehat{B}_{n}(g)-B(g;P)\right|_{p}\rightarrow 0.$
At the same time, by definition we have $V(\widetilde{g}_{P};P)-V(\widehat{g}_{\text{tradeoff}};P)\geq 0$. We thus conclude
$\sup_{P\in\mathcal{P}}\left|V(\widehat{g}_{\text{tradeoff}};P)-V(\widetilde{g}_{P};P)\right|\rightarrow_{p}0.$
Furthermore, we have $W(g_{P}^{\ast};P)\leq V(\widetilde{g}_{P};P)\leq W(\widetilde{g}_{P};P)$ as a by-product of Lemma 4.1 for any $P\in\mathcal{P}$. Putting these together, we have
$\inf_{P\in\mathcal{P}}\left\{W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)\right\}$
$\geq$ $\inf_{P\in\mathcal{P}}\left\{V(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)\right\}$
$\geq$ $\inf_{P\in\mathcal{P}}\left\{V(\widehat{g}_{\text{tradeoff}};P)-V(\widetilde{g}_{P};P)\right\}+\inf_{P\in\mathcal{P}}\left\{V(\widetilde{g}_{P};P)-W(g_{P}^{\ast};P)\right\}$
$\geq$ $\inf_{P\in\mathcal{P}}\left\{V(\widehat{g}_{\text{tradeoff}};P)-V(\widetilde{g}_{P};P)\right\}\rightarrow_{p}0.$
which proves uniform asymptotic welfare-efficiency.
By the uniform deviation bound shown above,applied to the difference between the value under $\widehat{g}_{\text{tradeoff}}$ and $g_{P}^{\ast}$, we have that uniformly with probability approaching one
$V(g_{P}^{\ast};P)-V(\widehat{g}_{\text{tradeoff}};P)\leq 0$
which is equivalent to $(B(\widehat{g}_{\text{tradeoff}};P)-k)_{+}\leq\frac{W(\widehat{g}_{\text{tradeoff}};P)-W(g_{P}^{\ast};P)}{\overline{\lambda}}$. ∎
Online Appendix to Empirical Welfare Maximization with Constraints
This Online Appendix contains proofs of supporting lemmas and additional results stated in the paper.
Appendix B Primitive assumptions and auxiliary lemmas
I first prove the optimal rule that solves the population constrained optimization problem takes the form of threshold. In Section B.2, I first provide primitive assumptions on the class of DGPs. I then prove in Lemma B.1, which establishes that these primitive assumptions imply Assumption 2.1.
In Section B.3, I verify Assumption 3.1 for settings where the observed sample comes from an RCT or an observational study, and the propensity score can be estimated efficiently based on parametric regressions.
B.1 Constrained optimal rule without functional form restriction
The population problem is to find rules based on $X_{i}$ that solves
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}E[\Gamma_{i}g(X_{i})]\ s.t.\ E[R_{i}g(X_{i})]<k$
By Law of Iterated Expectation, we can write the constrained optimization problem as
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}E[\gamma(X_{i})g(X_{i})]\ s.t.\ E[r(X_{i})g(X_{i})]$ $<k$
where $\gamma(X_{i})=E[\Gamma_{i}\mid X_{i}]$ and $r(X_{i})=E[R_{i}\mid X_{i}]$.
Claim B.1.
Let $d\mu=r(x)f(x)dx$ denote the positive measure. The constrained optimization problem is equivalent to
[…]
Let $X^{\ast}$ be the support of the solution $g^{\ast}$. It will take the form of $X^{\ast}=\{x:\frac{\gamma(x)}{r(x)}>c\}$ where $c$ is chosen so that $\mu(X^{\ast})=k$.
Proof.
Let $X$ be the support of any $g\neq g^{\ast}$ with $\mu(X)=k$. Then the objective function associated $g$ is
[…]
By definition of $X^{\ast}$, we have $\frac{\gamma(x)}{r(x)}>c$ for $x\in X^{\ast}\backslash X$ and $\frac{\gamma(x)}{r(x)}<c$ for $x\in X\backslash X^{\ast}$. Also note that $\mu$ is a positive measure. Then the above difference is lower bounded by
[…]
B.2 Primitive assumptions for contiguity
Assumption B.1.
[…]
Assumption B.2.
For all policies $g$, $B(g;P_{\theta})$ is twice continuously differentiable in $\theta$ at 00, and the derivatives are bounded from above and away from zero within an open neighborhood $\mathcal{N}_{\theta}$ of zero uniformly over $g\in\mathcal{G}$.
Lemma B.1.
Under Assumption B.1, the class $\mathcal{P}$ includes a sequence of data distribution $\{P_{h_{n}}\}$ that is contiguous to $P_{0}$ for every $h_{n}$ satisfying $\sqrt{n}h_{n}\rightarrow h$ e.g. take $h_{n}=h/\sqrt{n}$. This proves the first part of Assumption 2.1. Suppose further Assumption B.2 holds, then there exists some $h$ for the second part of Assumption 2.1 to hold,.
Proof.
Proof of Lemma B.1. By Theorem 7.2 of Vaart (1998), the log likelihood ratio process converges under $P_{0}$ (denoted with $\overset{p_{0}}{\rightsquigarrow}$) to a normal experiment
$\log\prod_{i=1}^{n}\frac{p_{h_{n}}}{p_{0}}(A_{i})$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h^{\prime}\dot{\ell}_{0}(A_{i})-\frac{1}{2}h^{\prime}I_{0}h+o_{P_{0}}(1)$ (25)
$\overset{p_{0}}{\rightsquigarrow}\mathcal{N}\left(-\frac{1}{2}h^{\prime}I_{0}h,h^{\prime}I_{0}h\right)$ (26)
where ˙$\dot{\ell}_{0}$ is the score and $I_{P_{0}}=E_{P_{0}}[\dot{\ell}_{0}(A_{i})\dot{\ell}_{0}(A_{i})^{\prime}]$ exists. The convergence in distribution of the log likelihood ratio to a normal with mean equal to $-\frac{1}{2}$ of its variance in (26) implies mutual contiguity $P_{0}^{n}\vartriangleleft\vartriangleright P_{h_{n}}^{n}$ by Le Cam’s first lemma (see Example 6.5 of Vaart (1998)). This proves the first part of the lemma.
By Taylor’s theorem with remainder we have for each policy $g$
$B(g;P_{h/\sqrt{n}})-B(g;P_{0})=\frac{h^{\prime}}{\sqrt{n}}\frac{\partial B(g;P_{0})}{\partial\theta}+\frac{1}{2}\frac{1}{n}h^{\prime}\frac{\partial^{2}B(g;P_{\tilde{\theta}_{n}})}{\partial\theta\partial\theta^{\prime}}h$
where $\tilde{\theta}_{n}$ is a sequence of values with $\tilde{\theta}_{n}\in[0,h/\sqrt{n}]$ that can depend on $g$. Take $h$ so that the first term is positive for policies with $B(g;P_{0})=k$. Such $h$ exists because we assume $\frac{\partial B(g;P_{0})}{\partial\theta}$ is bounded away from zero. For $g\in\mathcal{G}_{0}$ where $\mathcal{G}_{0}=\left\{g:B(g;P_{0})=k\right\}$, the constraints are violated under $P_{h_{n}}$ and furthermore (multiplying by $\sqrt{n}$)
$\sqrt{n}\cdot(B(g;P_{h_{n}})-k)$ $>h^{\prime}\frac{\partial B(g;P_{0})}{\partial\theta}>0$
for every $n$. This proves the second part of the lemma for $C=\inf_{g\in\mathcal{G}_{0}}\left|h^{\prime}\frac{\partial B(g;P_{0})}{\partial\theta}\right|$ ∎
B.3 Primitive assumptions and proofs for estimation quality
In this section, I verify that $\widehat{W}_{n}(g)$ and $\widehat{B}_{n}(g)$ satisfy Assumption 3.1 under primitive assumptions on the policy class and the OHIE.
Assumption B.3.
VC-class: The policy class $\mathcal{G}$ has a finite VC-dimension $v<\infty$.
To introduce the assumptions on the OHIE, I first recall the definitions of components in $\widehat{W}_{n}(g)$ and $\widehat{B}_{n}(g)$:
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i})$
for the doubly-robust scores
$\Gamma_{i}^{\ast}$ $=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$
$R_{i}^{\ast}$ $=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(V_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$
and observed characteristics $X_{i}$. Here $V_{i}$ collects the confounders in OHIE, namely household size (number of adults entered on the lottery sign-up form) and survey wave. Note that while $X_{i}$ can overlap with $V_{i}$, the policy $g(X_{i})$ needs not vary by $V_{i}$. In the OHIE example, the policy is based on number of children and income. However, conditional on household size and survey wave, income and number of children is independent of the lottery outcome in OHIE.
Recall that $\widehat{W}_{n}(g)$ and $\widehat{B}_{n}(g)$ are supposed to approximate net benefit $\Gamma$ and net excess cost $R$ of Medicaid eligibility. I provide more precise definitions for $\Gamma$ and $R$ as the primitive assumptions are stated in terms of their components.
Let $Y(1)$ be the (potential) subjective health when one is given Medicaid eligibility, and $Y(0)$ be the (potential) subjective health when one is not given Medicaid eligibility. Recall the definition for
$\Gamma=Y(1)-Y(0)$
We only observe the actual subjective health $Y_{i}$.
Let $M(1)$ be the (potential) enrollment in Medicaid when one is given Medicaid eligibility, and $M(0)$ be the (potential) enrollment in Medicaid when one is not given Medicaid eligibility.
Let $C(1)$ be the (potential) cost to the government when one is given Medicaid eligibility, and $C(0)$ be the (potential) cost to the government when one is not given Medicaid eligibility. Note that $C(0)=0$ by construction. Even when given eligibility, one might not enroll and thus incur zero cost to the government. So given the current expenditure is \$6,000 per enrollee, the implied expenditure per capita under eligibility policy $g(X)$ is actually \$$\$6,000\cdot E\left[M(1)g(X)\right]$. The reason is that the expected enrollment rate is only $Pr\{M(1)=1\mid X\in g\}$. So the per capita excess cost of Medicaid eligibility policy $g(X)$ relative to the current level is
$R=C(1)-\$6,000\cdot M(1).$
We only observe the actual cost ($C_{i}=D_{i}C_{i}(1)+(1-D_{i})C_{i}(0)$) and the actual Medicaid enrollment ($M_{i}$) in OHIE. We calculate $Z_{i}=C_{i}-\$6,000\cdot M_{i}$.
Assumption B.4.
Suppose for all $P\in\mathcal{P}$, the following statements hold for the OHIE:
Independent characteristics: $Pr\{D_{i}=1\mid V_{i},X_{i}\}=Pr\{D_{i}=1\mid V_{i}\}$
Unconfoundedness: $(Y(1),Y(0),C(1),M(1))\perp D_{i}|V_{i}$.
Bounded attributes: the support of variables $X_{i}$, $Y_{i}$ and $Z_{i}$ are bounded.
Strict overlap: There exist $\kappa\in(0,1/2)$ such that the propensity score satisfies $p(v)\in[\kappa,1-\kappa]$ for all $v\in\mathcal{V}$.
B.3.1 Uniform convergence of $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$
We want to show the recentered empirical processes $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$ converge to mean-zero Gaussian processes $G_{P}^{W}$ and $G_{P}^{B}$ with covariance functions $\Sigma_{P}^{W}(\cdot,\cdot)$ and $\Sigma_{P}^{B}(\cdot,\cdot)$ respectively uniformly over $P\in\mathcal{P}$. The covariance functions are uniformly bounded, with diagonal entries bounded away from zero uniformly over $g\in\mathcal{G}$. Take $\widehat{W}_{n}(\cdot)$ for example, the recentered empirical processes is
$\sqrt{n}\left(\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i})-E_{P}[\Gamma\cdot g(X_{i})]\right)$
and can be expressed as the sum of two terms
$\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})\cdot g(X_{i})+\sqrt{n}\left(\frac{1}{n}\sum_{i}\widetilde{\Gamma}_{i}\cdot g(X_{i})-E_{P}[\Gamma\cdot g(X_{i})]\right).$ (27)
Here $\widetilde{\Gamma}_{i}$ are the theoretical analogs
$\widetilde{\Gamma}_{i}$ $=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)+\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$
which is doubly-robust score with the theoretical propensity score and the CEF. A similar expansion holds for $\widehat{B}_{n}(\cdot)$ involving the theoretical analog
$\widetilde{R}_{i}=\gamma^{Z}(V_{i},1)+\frac{D_{i}}{p(V_{i})}\cdot\left(Z_{i}-\gamma^{Z}(V_{i},D_{i})\right).$
The following lemmas prove the uniform convergence of $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$.
The last part of the assumption is that we have a uniformly consistent estimate for the covariance function. I argue the sample analog
$\widehat{\Sigma}^{B}(g,g^{\prime})$ $=\frac{1}{n}\sum_{i}(\widetilde{R}_{i})^{2}\cdot g(X_{i})\cdot g^{\prime}(X_{i})-\left(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g(X_{i})\right)\cdot\left(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g^{\prime}(X_{i})\right)$
is a pointwise consistent estimate for $\widehat{\Sigma}^{B}(g,g^{\prime})$. To see this, note that the covariance function $\Sigma_{P}^{B}(\cdot,\cdot)$ maps $\mathcal{G}\times\mathcal{G}$ to $\mathbb{R}$. Under Assumption B.4 that $\mathcal{G}$ is a VC-class, the product $\mathcal{G}\times\mathcal{G}$ is also a VC-class (see e.g. Lemma 2.6.17 of van der Vaart and Wellner (1996)). By a similar argument leading to Lemma B.4 and B.6 , we conclude the uniform consistency of $\widehat{\Sigma}^{B}(\cdot,\cdot).$
Lemma B.2.
Let $\mathcal{G}$ be a VC-class of subsets of $\mathcal{X}$ with VC-dimension $v<\infty$. The following sets of functions from $\mathcal{A}$ to $\mathbb{R}$
$\mathcal{F}^{W}$ =$=\{\widetilde{\Gamma}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
$\mathcal{F}^{B}$ =$=\{\widetilde{R}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
are VC-subgraph class of functions with VC-dimension less than or equal to $v$ for all $P\in\mathcal{P}$. For notational simplicity, we suppress the dependence of $\mathcal{F}$ on $P$.
Lemma B.3.
For all $P$ in the family $\mathcal{P}$ of distributions satisfying Assumption B.4, for all $g\in\mathcal{G}$ we have
$E_{P}[\widetilde{\Gamma}_{i}\cdot g(X_{i})]$ $=W(g;P)$
$E_{P}[\widetilde{R}_{i}\cdot g(X_{i})]$ $=B(g;P)$
Lemma B.4.
Let $\mathcal{G}$ satisfy Assumption B.4 (VC). Let $U_{i}$ be a mean-zero bounded random vector of fixed dimension i.e. there exists $M<\infty$ such that i.e. $U_{i}\in[-M/2,M/2]$ almost surely under all $P\in\mathcal{P}$. Then the uniform deviation of the sample average of vanishes
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{n}\sum_{i}U_{i}\cdot g(X_{i})\right|$ $\rightarrow_{a.s.}\mathbf{0}.$
Lemma B.5.
Let $\mathcal{P}$ be a family of distributions satisfying Assumption B.4. Let $\mathcal{G}$ satisfy Assumption B.4 (VC). Then $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ are $P$-Donsker for all $P\in\mathcal{P}$. That is, the empirical process indexed by $g\in\mathcal{G}$
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{\Gamma}_{i}\cdot g(X_{i})-W(g;P))$
converge to a Gaussian process $\mathcal{GP}(0,\Sigma_{P}^{W}(\cdot,\cdot))$ uniformly in $P\in\mathcal{P}$, and the empirical process indexed by $g\in\mathcal{G}$
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g(X_{i})-B(g;P))$
converge to a Gaussian process $\mathcal{GP}(0,\Sigma_{P}^{B}(\cdot,\cdot))$ uniformly in $P\in\mathcal{P}$.
Lemma B.6.
Let $\mathcal{P}$ be a family of distributions satisfying Assumption B.4. Let $\mathcal{G}$ satisfy Assumption B.4 (VC). Then the estimation errors vanishes
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(R_{i}^{\ast}-\widetilde{R}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
B.3.2 Proofs of auxiliary lemmas
Proof.
Proof of Lemma B.2. This lemma follows directly from Lemma A.1 of Kitagawa and Tetenov (2018). ∎
Proof.
Proof of Lemma B.3. Under Assumption B.4, we prove each $\widetilde{\Gamma}_{i}$ is an conditionally unbiased estimate for $\Gamma$:
$E_{P}[\widetilde{\Gamma}_{i}g(X_{i})]=E_{P}[E_{P}[\widetilde{\Gamma}_{i}\mid X_{i}]g(X_{i})]=E_{P}[E_{P}[\Gamma\mid X_{i}]g(X_{i})]=W(g;P).$
We focus on $E_{P}[\widetilde{\Gamma}_{i}\mid X_{i}]=E_{P}[E_{P}[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]\mid X_{i}]$. However, conditional on $V_{i}$, by unconfoundedness and strict overlap we have
$E_{P}[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]$ $=E[Y_{i}\mid V_{i},X_{i},D_{i}=1]-E[Y_{i}\mid V_{i},X_{i},D_{i}=0]$
$=E[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}]=E[\Gamma_{i}\mid V_{i},X_{i}]$
Specifically
$E[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]$ $=E\left[\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)\mid V_{i},X_{i}\right]+E\left[\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)\mid V_{i},X_{i}\right]$
$=E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right]+$
$E\left[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}\right]-(E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right])$
$=E[\Gamma\mid V_{i},X_{i}]$
Specifically, we expand $E[\alpha(V_{i},D_{i})Y_{i}\mid V_{i},X_{i}]$
$=E\left[\frac{Y_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot Pr\{D_{i}=1\mid V_{i},X_{i}\}-E\left[\frac{Y_{i}}{1-p(V_{i})}\mid D_{i}=0,X_{i}\right]\cdot Pr\{D_{i}=0\mid V_{i},X_{i}\}$
$=E\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=1,V_{i},X_{i}\right]-E\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=0,V_{i},X_{i}\right]$
$=E\left[Y_{i}(1)\mid V_{i},X_{i}\right]-E\left[Y_{i}(0)\mid V_{i},X_{i}\right]$
where the second line holds by independent characteristic such that
$Pr\{D_{i}=1\mid V_{i},X_{i}\}=Pr\{D_{i}=1\mid V_{i}\}\eqqcolon p(V_{i})$
and the last line follows from unconfoundedness. Similarly, we show
$E\left[\alpha(V_{i},D_{i})\cdot\gamma^{Y}(V_{i},D_{i})\mid V_{i},X_{i}\right]=E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right]$
Similar argument holds for $E_{P}[\widetilde{R}_{i}\cdot g(X_{i})]$, with the only modification:
$E[\frac{D_{i}}{p(V_{i})}C_{i}\mid V_{i},X_{i}]=E\left[\frac{C_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot Pr\{D_{i}=1\mid V_{i},X_{i}\}$
$=E[C_{i}(1)\mid V_{i},X_{i}]$
Proof.
Proof of Lemma B.4. Denote the following set of functions from $\mathcal{U}$ to $\mathbb{R}$
$\mathcal{F}^{U}$ =$=\{U_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
and it has uniform envelope $\bar{F}=M/2$ since $U_{i}$ is bounded. This envelop function is bounded uniformly over $\mathcal{P}$. Also, by Assumption B.4 (VC) and Lemma B.2, $\mathcal{F}^{U}$ is VC-subgraph class of functions with VC-dimension at most $v$. By Lemma 4.14 and Proposition 4.18 of Wainwright (2019), we conclude that $\mathcal{F}^{U}$ has Rademacher complexity $2\sqrt{M^{2}\frac{v}{n}}$. Then by Proposition 4.12 of Wainwright (2019) we conclude that $\mathcal{F}^{U}$ are $P$-Glivenko–Cantelli for each $P\in\mathcal{P}$, with an $O(\sqrt{\frac{v}{n}})$ rate of convergence. Note that this argument does not use any constants that depend on $P$ but only $M$ and $v$, so we can actually get uniform convergence over $\mathcal{P}$. ∎
Proof.
Proof of Lemma B.5. Note that Assumption B.4 imply that $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ have uniform envelope $\bar{F}=M/(2\kappa)$. $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ thus have square integrable envelop functions uniformly over $\mathcal{P}$. Also, by Assumption B.4 (VC) and Lemma B.2, $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ are VC-subgraph class of functions with VC-dimension at most $v$. Even though both $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ depend on $P$, a similar argument for Theorem 1 in Rai (2019) show that $\mathcal{F}^{W}$ and $\mathcal{F}^{B}$ are $P$-Donsker uniformly in $P\in\mathcal{P}$. ∎
Proof.
Proof of Lemma B.6. We focus on the deviation in $\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i}$. The deviation in $R_{i}^{\ast}-\widetilde{R}_{i}$ can be proven to vanish in a similar manner. Denote $\Delta\gamma^{Y}(V_{i})=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)$. For any fixed policy $g$, we expand the deviation $\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})g(X_{i})$ into three terms
$\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(\widehat{\alpha}(V_{i},D_{i})-\alpha(V_{i},D_{i})\right)\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$ (28)
$+\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(\Delta\widehat{\gamma}^{Y}(V_{i})-\Delta\gamma^{Y}(V_{i})-\alpha(V_{i},D_{i})\cdot\left(\widehat{\gamma}^{Y}(V_{i},D_{i})-\gamma^{Y}(V_{i},D_{i})\right)\right)$
$-\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(\widehat{\gamma}^{Y}(V_{i},D_{i})-\gamma^{Y}(V_{i},D_{i})\right)\cdot\left(\widehat{\alpha}(V_{i},D_{i})-\alpha(V_{i},D_{i})\right)\cdot$
Denote these three summands by $D_{1}(g)$, $D_{2}(g)$ and $D_{3}(g)$. We will bound all three summands separately. Recall we use the full sample to estimate the propensity score and the CEF with a saturated model. The purpose of the above expansion is to separately bound the estimation error from the estimated CEF and propensity score, and the deviation from taking sample averages. For cross-fitted estimators for the propensity score and the CEF, a similar bound can be found in Athey and Wager (2021).
Uniform consistency of the estimated CEF and propensity score
Denote with $b(V_{i},D_{i})$ the dictionary that spans $(V_{i},D_{i})$, and $b(V_{i},)$ the dictionary that spans $V_{i}$. The saturated models are therefore parameterized as $\gamma^{Z}(V_{i},D_{i})=\gamma^{\prime}b(V_{i},D_{i})$ and $p(V_{i})=\beta^{\prime}b(V_{i})$. Under standard argument, the OLS estimators $\widehat{\gamma}$ and $\widehat{\beta}$ are asymptotically normal uniformly over $P\in\mathcal{P}$:
$\sqrt{n}\cdot\left(\widehat{\gamma}-\gamma\right)=O_{P}(1),\ \sqrt{n}\cdot\left(\widehat{\beta}-\beta\right)=O_{P}(1).$
Furthermore, the in-sample $L_{2}$ errors from the estimated CEF and propensity score vanish. Consider
$\frac{1}{n}\sum_{i}\left(\widehat{\gamma}^{Z}(V_{i},D_{i})-\gamma^{Z}(V_{i},D_{i})\right)^{2}=\frac{1}{n}\sum_{i}\left(\left(\widehat{\gamma}-\gamma\right)^{\prime}b(V_{i},D_{i})\right)^{2}=\left(\widehat{\gamma}-\gamma\right)^{\prime}\widehat{M}\left(\widehat{\gamma}-\gamma\right)$
where $\widehat{M}=\frac{1}{n}\sum_{i}b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}.$ It converges in probability to a fixed matrix $M=E[b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}]$. So the in-sample $L_{2}$ error from the estimated CEF vanishes at the rate of $n^{-1/2}$.
Similarly, consider expanding $\frac{1}{n}\sum_{i}\left(\widehat{\alpha}(V_{i},D_{i})-\alpha(V_{i},D_{i})\right)^{2}$ as
$\frac{1}{n}\sum_{i}\left(\frac{1}{\widehat{\beta}^{\prime}b(V_{i})}-\frac{1}{\beta^{\prime}b(V_{i})}\right)^{2}D_{i}^{2}+\left(\frac{1}{1-\widehat{\beta}^{\prime}b(V_{i})}-\frac{1}{1-\beta^{\prime}b(V_{i})}\right)^{2}\left(1-D_{i}\right)^{2}$
With a first-order Taylor approximation, for each term in the summand, the dominating term would be
$\frac{1}{n}\sum_{i}\left(\left(\widehat{\beta}-\beta\right)^{\prime}\frac{-b(V_{i})}{(\beta^{\prime}b(V_{i}))^{2}}\right)^{2}D_{i}^{2}$ $=\left(\widehat{\beta}-\beta\right)^{\prime}\left(\frac{1}{n}\sum_{i}\frac{b(V_{i})b(V_{i})^{\prime}}{(\beta^{\prime}b(V_{i}))^{2}}D_{i}^{2}\right)\left(\widehat{\beta}-\beta\right)$
where the middle term converges to a fixed matrix as implied by $\beta^{\prime}b(V_{i})$ being bounded away from zero and one. So the in-sample $L_{2}$ error from the estimated propensity score also vanishes at the rate of $n^{-1/2}$.
Bounding the deviation
We now bound each term in (28). Plugging in the first-order Taylor approximation with a remainder term to the estimated propensity score, we have
$D_{1}(g)=$ $\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)\cdot$
$\left(\left(\frac{1}{\widehat{\beta}^{\prime}b(V_{i})}-\frac{1}{\beta^{\prime}b(V_{i})}\right)D_{i}+\left(\frac{1}{1-\widehat{\beta}^{\prime}b(V_{i})}-\frac{1}{1-\beta^{\prime}b(V_{i})}\right)\left(1-D_{i}\right)\right)$
$=$ $\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(\widehat{\beta}-\beta\right)^{\prime}\frac{-b(V_{i})}{(\beta^{\prime}b(V_{i}))^{2}}\cdot D_{i}\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)+$
$\frac{1}{\sqrt{n}}\sum_{i}g(X_{i})\left(\widehat{\beta}-\beta\right)^{\prime}\frac{b(V_{i})b(V_{i})^{\prime}}{(\widetilde{\beta}^{\prime}V_{i})^{3}}\left(\widehat{\beta}-\beta\right)\cdot D_{i}\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$
$=$ $\underbrace{\sqrt{n}\left(\widehat{\beta}-\beta\right)}_{O_{P}(1)}^{\prime}\underbrace{\frac{1}{n}\sum_{i}g(X_{i})\frac{-b(V_{i})}{(\beta^{\prime}b(V_{i}))^{2}}\cdot D_{i}\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)}_{o_{P}(1)}+o_{P}(1)$
where $\widetilde{\beta}$ is a sequence between $\widehat{\beta}$ and $\beta$. This remainder term therefore converges to zero. Uniform convergence of the sample average follows from Lemma B.4: the random vector $\frac{-b(V_{i})}{(\beta^{\prime}b(V_{i}))^{2}}\cdot D_{i}\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$ is mean-zero, has fixed dimension, and bounded.
We can decompose $D_{2}(g)$ into the product of two terms
$D_{2}(g)=\underbrace{\sqrt{n}\left(\widehat{\gamma}-\gamma\right)}_{O_{p}(1)}^{\prime}\frac{1}{n}\sum_{i}g(X_{i})\left(\Delta b(V_{i},D_{i})-\alpha(V_{i},D_{i})\cdot b(V_{i},D_{i})\right)$
Uniform convergence of the sample average again follows from Lemma B.4. We thus conclude $D_{1}(g)$ and $D_{2}(g)$ vanish uniformly over $g\in\mathcal{G}$ and over $P\in\mathcal{P}$.
For $D_{3}(g)$, we apply the Cauchy-Schwarz inequality to note that
$D_{3}(g)\leq\sqrt{n}\cdot\sqrt{\frac{1}{n}\sum_{i}\left(\widehat{\gamma}^{Y}(V_{i},D_{i})-\gamma^{Y}(V_{i},D_{i})\right)^{2}}\cdot\sqrt{\frac{1}{n}\sum_{i}\left(\widehat{\alpha}(V_{i},D_{i})-\alpha(V_{i},D_{i})\right)^{2}}$
The terms in the square root are the in-sample $L_{2}$ errors from the estimated CEF and propensity score, which vanish at the rate of $n^{-1/2}$ uniformly over $P\in\mathcal{P}$ as shown in the paragraph above. We thus conclude $D_{3}(g)$ vanishes uniformly over $g\in\mathcal{G}$ and over $P\in\mathcal{P}$. ∎
Appendix C Computation
The policy class considered in this paper $g(x)=\mathbf{1}\{\beta^{\prime}x\geq 0\}$ consists of linear eligibility score. Therefore I follow (Kitagawa and Tetenov, 2018, Online Appendix C) to set up the problems of sample-analog rule (9) and trade-off rule (18) as a Mixed Integer Linear Programming (MILP), which is more efficient than solving them through a search over $x$. Due to the estimated constraint, I add linear constraints to the version of Kitagawa and Tetenov (2018) and instead solve for the trade-off rule in two parts:
constrained unconstrained (29)
$\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\Gamma_{i}^{\ast}z_{i}$ $\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\Gamma_{i}^{\ast}z_{i}+\overline{\lambda}(\sum_{i=1}^{n}R_{i}^{\ast}z_{i}-n\cdot k)$
$s.t.\ \frac{X_{i}^{\prime}\beta}{C_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{C_{i}},$ $s.t.\ \frac{X_{i}^{\prime}\beta}{C_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{C_{i}},$
$z_{i}\in\{0,1\},$ $z_{i}\in\{0,1\},$
$\sum_{i=1}^{n}R_{i}^{\ast}z_{i}<n\cdot k.$ $\sum_{i=1}^{n}R_{i}^{\ast}z_{i}\geq n\cdot k.$
where constants $C_{i}=\sup_{\beta\in B}|X_{i}^{\prime}\beta|$. The additional binary parameters $(z_{1},\dots,z_{n})$ replace the policy functions $g(X_{i})=\mathbf{1}\{\beta^{\prime}X_{i}\geq 0\}$. The sample-analog rule $\widehat{g}_{\text{sample}}$ is the constrained solution on the left hand side. The trade-off rule $\widehat{g}_{\text{tradeoff}}$ is the better solution across constrained and unconstrained policies in (29).
To modify the sample-analog rule to control the probability of selecting infeasible policies by $\widehat{g}_{\alpha}$ as proposed in Section 3.3, I introduce quadratic constraints to (29). First, for a conventional level of $\alpha$, constructing the tightened constraint $\hat{\mathcal{G}}_{\alpha}$ requires an estimate for the critical value $c_{\alpha}$, the $\alpha\%$-quantile from $\underset{g\in\mathcal{G}}{\inf}\ \frac{G_{P}^{B}(g)}{\Sigma_{P}^{B}(g,g)^{1/2}}$, the infimum of the Gaussian process $G_{P}^{B}(\cdot)\sim\mathcal{GP}(0,\Sigma_{P}^{B}(\cdot,\cdot))$. In practice, I construct a grid on $\mathcal{G}$ as
$\mathcal{\widetilde{G}}=\left\{g(x)=\mathbf{1}\{income\cdot 1\{numchild=j\}\leq y_{j}\}:j\in\{0,1,\geq 2\},y_{j}\in\{0,50,100,\dots,500\}\right\}$ (30)
for characteristics $x=(income,\ numchild)$. This grid thus consists of income thresholds every 50% of the federal poverty level, and the thresholds can vary with number of children. I then approximate the infimum over infinite-dimensional $\mathcal{G}$ by the minimum over $\mathcal{\widetilde{G}}$ with estimated covariance i.e. $\underset{g\in\widetilde{\mathcal{G}}}{\min}\ \frac{\widetilde{h}(g)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}$. Here $\widetilde{h}(\cdot)\sim\mathcal{\mathcal{N}}(0,\widehat{\Sigma}^{B})$ is a Gaussian vector indexed by $g\in\widetilde{\mathcal{G}}$, with $\widehat{\Sigma}^{B}$ is sub-matrix of the covariance estimate $\widehat{\Sigma}^{B}(\cdot,\cdot)$ for $g\in\widetilde{\mathcal{G}}$. Based on 10,000 simulation draws I estimate $c_{\alpha}$ to be -2.56 for $\alpha=5\%$. The validity of this approximation is given by the uniform consistency of the covariance estimator under Assumption 3.1.
Second, to select policies in $\hat{\mathcal{G}}_{\alpha}$ is equivalent to invert a test: $\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}$. This test inversion can be written as two constraint (note that $c_{\alpha}<0$). The first constraint is linear: $\sum_{i=1}^{n}R_{i}^{\ast}z_{i}\leq n\cdot k$ and the second is $n\cdot\left(\widehat{B}_{n}(g)-k\right)^{2}>c_{\alpha}^{2}\cdot\widehat{\Sigma}^{B}(g,g)$ which translates to a quadratic constraint
$n\left(\frac{\sum_{i=1}^{n}R_{i}^{\ast}z_{i}}{n}-k\right)^{2}>c_{\alpha}^{2}\left(\left(\frac{1}{n}\sum_{i=1}^{n}(R_{i}^{\ast})^{2}z_{i}\right)-\left(\frac{\sum_{i=1}^{n}R_{i}^{\ast}z_{i}}{n}\right)^{2}\right)$
$\Rightarrow$ $\frac{c_{\alpha}^{2}}{n}\sum_{i=1}^{n}(R_{i}^{\ast})^{2}z_{i}+2k\sum_{i=1}^{n}R_{i}^{\ast}z_{i}-\left(\frac{1}{n}+\frac{c_{\alpha}^{2}}{n^{2}}\right)(\sum_{i=1}^{n}R_{i}^{\ast}z_{i})^{2}<nk^{2}$